### v2

# Endoscopic character identities for depth-zero supercuspidal representations of $\mathrm{SL}_{2}$

Abstract.
We consider the depth-zero supercuspidal $L$-packets of $\mathrm{SL}_{2}$. With the aid of the classical character formulas of Sally-Shalika, we prove the endoscopic character identities. We comment on the truth-status of Theorem 15.2 in the 2011 article by Adler, DeBacker, Sally and Spice.
Key words and phrases:
Endoscopy, characters, $p$-adic, $\mathrm{SL}_{2}$
1. Introduction
We consider the depth-zero supercuspidal $L$-packets of $\mathrm{SL}_{2}$. With the aid of the classical character formulas of Sally-Shalika, we prove the endoscopic character identities.
Let $F$ be a non-archimedean local field of zero characteristic (that is, $F$ is a finite extension of $\mathbb{Q}_{p}$). We shall assume that $p>2$. Let $q=q_{F}$ denote the cardinality of the residue field of $F$. Let $\mathrm{v}(x)$ denote the valuation of $x\in F^{\times}$. Let $\mathfrak{o}_{F}$ denote the ring of integers of $F$, let $\varpi$ be a uniformiser for $F$, let $\mathfrak{p}_{F}=\varpi\mathfrak{o}_{F}$ be the maximal ideal of $\mathfrak{o}_{F}$, and let $\varepsilon$ be a fixed nonsquare element in $\mathfrak{o}_{F}^{\times}$. In that case, $\{1,\varpi,\varepsilon,\varepsilon\varpi\}$ is a set of representatives of $F^{\times}/(F^{\times})^{2}$.
Let $G=\mathrm{SL}_{2}(F)$, and let $\mathrm{Z}(G)$ denote the center of $G$. We shall focus on the elliptic torus
(1.1) $T^{\varepsilon}:=\left\{\left(\begin{matrix}a&b\cr b\varepsilon&a\end{matrix}\right)\;:\;a,b\in\mathfrak{o}_{F},a^{2}-\varepsilon b^{2}=1\right\}.$
This elliptic torus $T^{\varepsilon}$ is a representative of the single stable conjugacy class of unramified elliptic maximal $F$-tori in $\mathrm{SL}_{2}(F)$.
We will define $f$ as follows:
(1.2) $f\colon T^{\varepsilon}\to\mathbb{Z},\quad\quad\left(\begin{matrix}a&b\cr b\varepsilon&a\end{matrix}\right)\mapsto(-q)^{\mathrm{v}(b)}.$
We denote by $T^{\varepsilon}_{1}$ the first compact open subgroup in the Moy-Prasad filtration of $T^{\varepsilon}$, namely
(1.3) $T^{\varepsilon}_{1}:=\left\{\left(\begin{matrix}a&b\cr b\varepsilon&a\end{matrix}\right)\in T^{\varepsilon}\;:\;a\in 1+\mathfrak{p}_{F},\;b\in\mathfrak{p}_{F}\;\right\}.$
Let $\mathcal{B}(G)$ be the Bruhat-Tits building of $G$. For $x\in\mathcal{B}(G)$, we denote by $G_{x}$ the associated parahoric subgroup and by $G_{x,0+}$ the pro $p$ unipotent radical of $G_{x}$. We set
(1.4) $G_{0+}:=\bigcup_{x\in\mathcal{B}(G)}G_{x,0+}.$
We will say, following [ADSS], that $\gamma$ is near the identity if $\gamma\in G_{0+}$, and far from the identity if $\gamma\notin\mathrm{Z}(G)G_{0+}$. If $\gamma$ is a regular semisimple element in $T^{\varepsilon}$ which is near the identity, then we have $\gamma\in T_{1}^{\varepsilon}$, see §2.0.4.
We fix a Whittaker datum $\mathfrak{w}$. Let $\phi$ be a Langlands parameter for a depth-zero supercuspidal $L$-packet of $\mathrm{SL}_{2}(F)$ ($\phi$ is uniquely determined up-to conjugation by $\mathrm{PGL}_{2}(\mathbb{C})$), and let $\Pi_{\phi}(\mathrm{SL}_{2}(F))$ denote the $L$-packet of $\phi$. We will specify which element in $\Pi_{\phi}(\mathrm{SL}_{2}(F))$ is $\mathfrak{w}$-generic.
Let $S_{\phi}$ denote the component group of the centralizer in $\mathrm{PGL}_{2}(\mathbb{C})$ of the image of $\phi$, and $\mathrm{Irr}(S_{\phi})$ the set of isomorphism classes of irreducible representations of $S_{\phi}$. We have a bijective map
(1.5) $\iota_{\phi}\colon\Pi_{\phi}(\mathrm{SL}_{2}(F))\to\mathrm{Irr}(S_{\phi})$
that sends the $\mathfrak{w}$-generic representation to the trivial character.
Let $s$ be a semisimple element in $S_{\phi}$. We attach to $(\phi,s)$ the virtual character
(1.6) $\Theta_{\phi,s}:=\sum_{\pi\in\Pi_{\phi}(\mathrm{SL}_{2}(F))}\mathrm{trace}(\iota_{\phi}(\pi))(s)\cdot\Theta_{\pi}.$
Let $\widehat{H}$ denote the identity component of the centralizer of $s$ in $\widehat{G}$. Then $\phi$ will factorize as follows:
[…]
The stable character $S\Theta_{\phi^{H}}$ is defined as
$S\Theta_{\phi^{H}}:=\Theta_{\phi^{H},1}.$
We set $\mathcal{H}:=\hat{H}\cdot\phi(\mathbf{W}_{F})$ and denote by $\mathfrak{e}=(H,\mathcal{H},s,\xi)$ the corresponding endoscopic datum, where $\xi\colon\mathcal{H}\hookrightarrow\hat{G}$. Then the endoscopic character identity for $\mathrm{SL}_{2}$ is
(1.7) $\Theta_{\phi,s}(\gamma)=\sum_{\gamma^{H}}\Delta[\mathfrak{w},\mathfrak{e}](\gamma^{H},\gamma)S\Theta_{\phi^{H}}(\gamma^{H}),$
where $\gamma$ and $\gamma^{H}$ are related, and $\Delta[\mathfrak{w},\mathfrak{e}]$ is the Langlands-Shelstad transfer factor.
The endoscopic character identities, for the depth-zero supercuspidal $L$-packets of $\mathrm{SL}_{2}$, are stated in Theorems 1.8 and 1.10. With the aid of the classical Sally-Shalika character formulas, we provide explicit proofs. We call an $L$-parameter $\phi$ regular (resp. non-singular) if the representations inthe $L$-packet $\Pi_{\phi}(\mathrm{SL}_{2}(F))$ are regular (resp. non-singular) in the sense of [Ka2, Ka3]. We note that the regular $L$-parameter in Theorem 1.8 fits into the framework of [Ka2], whereas the non-regular $L$-parameter in Theorem 1.10 does not (but, being non-singular, fits into the more general framework of [Ka3]).
The characters of the discrete series for $\mathrm{SL}_{2}(F)$ were obtained by Sally-Shalika [SS]. In particular, for the depth-zero supercuspidal $L$-packets, the characters appear in Table 3 of [SS]. They are given as locally integrable functions on the maximal torus of $\mathrm{SL}_{2}(F)$ and on the elliptic tori in $\mathrm{SL}_{2}(F)$. Many years later, proofs were supplied by Adler, DeBacker, Sally and Spice in [ADSS]. For the most part, we find it more convenient to refer to this paper [ADSS].
Theorem 1.8.
Let $\gamma$ be a regular, semisimple element in $T^{\varepsilon}$. Let $\Pi_{\phi}(\mathrm{SL}_{2}(F))$ be the depth-zero supercuspidal $L$-packet with regular $L$-parameter $\phi$. Let $H$ be the unramified elliptic torus $T^{\varepsilon}$. Then we have
(1.9) $\sum_{\gamma^{H}}\Delta[\mathfrak{w},\mathfrak{e}](\gamma^{H},\gamma)S\Theta_{\phi^{H}}(\gamma)$ $=$ $-f(\gamma)(\psi(\gamma)+\psi(1/\gamma)).$
From the Sally-Shalika character formulas we have
$\Theta_{\phi,s}(\gamma)$ $=$ $-\psi(\gamma)-\psi(1/\gamma)\quad\textrm{far from the identity}$
$\Theta_{\phi,s}(\gamma)$ $=$ $-2f(\gamma)\quad\textrm{near the identity}$
where $s$ is the non-trivial element of $S_{\phi}$, the elements $\gamma$ and $\gamma^{H}$ are related, and $\psi$ is a character of $T^{\varepsilon}$ such that $\psi\neq\psi^{-1}$.
Far from the identity, the right-hand-side of (1.9) is
$-\psi(\gamma)-\psi(1/\gamma)$
since $f(\gamma)=1$ by Lemma 2.18. Near the identity, the right-hand-side of (1.9) is
$-2f(\gamma)$
since $\gamma$ and $1/\gamma$ are both near the identity and $\psi$ has depth zero. This creates consistency in the formulas in Theorem 1.8, and mirrors the calculations in [ADSS]: in that paper, the character formulas are done separately, according as $\gamma$ is far from the identity or near the identity. Note that Theorem 1.8 fits into the framework of [Ka2] since $\phi$ is a regular parameter.
Theorem 1.10.
Let $\gamma$ be a regular, semisimple element in $T^{\varepsilon}$. Let $\Pi_{\phi}$ be the unique depth-zero supercuspidal $L$-packet with non-regular parameter $\phi$. Let $\psi_{0}$ be the unique quadratic character of $T^{\varepsilon}$. Then
(1.11) $\sum_{\gamma^{H}}\Delta[\mathfrak{w},\mathfrak{e}](\gamma^{H},\gamma)S\Theta_{\phi^{H}}(\gamma^{H})$ $=$ $-2f(\gamma)\psi_{0}(\gamma).$
From the Sally-Shalika character formulas, far from the identity, we have
$\Theta_{\phi,s}(\gamma)=-2\psi_{0}(\gamma)$
Near the identity, we have
$\Theta_{\phi,s}(\gamma)=-2f(\gamma).$
On Theorem 15.2 in [ADSS]. This theorem is unreliable: it is incompatible with Proposition 13.14 in [ADSS] and with the endoscopic character identity, see §5. We note that Theorem 15.2 occurs already in the 1968 article [SS] of Sally-Shalika, see the 7th (incorrect) formula in Table 3 of [SS]. We have used Proposition 13.14 in [ADSS] instead.
On the issue of uniqueness. In the context of a regular parameter $\phi$, the uniqueness of the map $\iota_{\phi}\colon\Pi_{\phi}\to\mathrm{Irr}(S_{\phi})$ is clear. For the $\mathfrak{w}$-generic element in $\Pi_{\phi}$ is sent to the trivial character of $S_{\phi}$, whereas the other element in $\Pi_{\phi}$ is sent to the non-trivial character of $S_{\phi}$.
In the context of the non-regular parameter $\phi$, the left-hand-side of (1.7) depends, in its definition, on the map $\iota_{\phi}\colon\Pi_{\phi}\to\mathrm{Irr}(S_{\phi})$. The unique map $\iota_{\phi}:\Pi_{\phi}\to\mathrm{Irr}(S_{\phi})$ for which (1.7) is true is defined in §4.
Acknowledgement. We would like to thank Loren Spice for a valuable exchange of emails and Tasho Kaletha for sending us, on more than one occasion, constructive criticism.
2. Unramified elliptic maximal $F$-tori
2.0.1. Elliptic tori
There are two tori relevant for us: Let $\varepsilon$ be a fixed nonsquare element in $\mathfrak{o}_{F}^{\times}$, then
(2.1) $T^{\varepsilon}:=\left\{\left(\begin{matrix}a&b\cr b\varepsilon&a\end{matrix}\right)\;:\;a,b\in\mathfrak{o}_{F},a^{2}-\varepsilon b^{2}=1\right\}$
and
(2.2) $T^{\varepsilon,\varpi}:=\left\{\left(\begin{matrix}a&b\varpi^{-1}\cr b\varepsilon\varpi&a\end{matrix}\right)\;:\;a,b\in\mathfrak{o}_{F},a^{2}-\varepsilon b^{2}=1\right\}.$
These two tori admit the same splitting field, namely $F(\sqrt{\varepsilon})$ and are therefore stably conjugate, or, equivalently, $\mathrm{GL}_{2}(F)$-conjugate. The tori $T^{\varepsilon}$ and $T^{\varepsilon,\varpi}$ are $\mathrm{GL}_{2}(F)$-conjugate, thanks to the following identity in $\mathrm{GL}_{2}(F)$:
(2.3) $\left(\begin{array}[]{cc}1&0\\ 0&\varpi\end{array}\right)\left(\begin{array}[]{cc}a&b\\ \varepsilon b&a\end{array}\right)\left(\begin{array}[]{cc}1&0\\ 0&\varpi^{-1}\end{array}\right)=\left(\begin{array}[]{cc}a&\varpi^{-1}b\\ \varepsilon\varpi b&a\end{array}\right).$
There is a single stable conjugacy class of unramified elliptic maximal $F$-tori, represented by $T^{\varepsilon}$. It splits into two distinct $G$-conjugacy classes, represented by $T^{\varepsilon}$ and $T^{\varepsilon,\varpi}$.
Let $E^{1}$ denote the group of norm $1$ elements in $E$. We have, following [ADSS, §3.1], canonical isomorphisms of $F$-tori:
$T^{\varepsilon}\to E^{1},\quad\left(\begin{matrix}a&b\cr b\varepsilon&a\end{matrix}\right)\mapsto a+b\sqrt{\varepsilon}$
$T^{\varepsilon,\varpi}\to E^{1},\quad\left(\begin{matrix}a&b\varpi^{-1}\cr b\varepsilon\varpi&a\end{matrix}\right)\mapsto a+b\sqrt{\varepsilon}.$
Given $a+b\sqrt{\varepsilon}\in E^{1}$ we shall, following [ADSS, §3.1] write
$\mathrm{Im}_{\varepsilon}(a+b\sqrt{\varepsilon})=b.$
We have
(2.4) $\mathrm{Im}_{\varepsilon}(\mathbf{g}\cdot\gamma)=\mathrm{Im}_{\varepsilon}(\gamma),$
for every $\gamma\in T^{\varepsilon}$, where
(2.5) $\mathbf{g}:=\left(\begin{array}[]{cc}1&0\\ 0&\varpi\end{array}\right)\in\mathrm{GL}_{2}(F).$
The equation (2.3) can be rephrased as the following statement:
(2.6) $\mathbf{g}\cdot T^{\varepsilon}=\mathbf{g}T^{\varepsilon}\mathbf{g}^{-1}=T^{\varepsilon,\varpi}.$
We also note that
$\mathrm{sgn}_{\varepsilon}(x)=(-1)^{\mathrm{v}(x)}$
for all $x\in F$, see (1.2) in [ADSS]. Then we have, using (2.4),
(2.7) $\mathrm{sgn}_{\varepsilon}(\varpi^{-1}\mathrm{Im}_{\varepsilon}(\mathbf{g}\cdot\gamma)$ $=$ $\mathrm{sgn}_{\varepsilon}(\varpi^{-1}\mathrm{Im}_{\varepsilon}(\gamma))$
(2.8) $=$ $\mathrm{sgn}_{\varepsilon}(\varpi^{-1})\mathrm{sgn}_{\varepsilon}(\mathrm{Im}_{\varepsilon}(\gamma))$
(2.9) $=$ $(-1)^{\mathrm{v}(\varpi^{-1})}\mathrm{sgn}_{\varepsilon}(\mathrm{Im}_{\varepsilon}(\gamma))$
(2.10) $=$ $-\mathrm{sgn}_{\varepsilon}(\mathrm{Im}_{\varepsilon}(\gamma)).$
2.0.2. Weyl discriminant
In [ADSS, Definition 3.7], the Weyl discriminant $D_{G}$ is defined as follows:
$D_{G}\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right)=(a+d)^{2}-4.$
The usual definition of the Weyl discriminant $D^{G}$ is the following. If $\gamma\in T(F)$ then
(2.11) $D^{G}(\gamma)=|\Pi_{\alpha}(\alpha(\gamma)-1)|^{1/2}$
where the product is over all roots of $T$ in $G$.
Let $A$ denote the diagonal subgroup of $\mathrm{SL}_{2}(F)$ and let $A_{0}$ denote the compact elements in $A$. According to [ADSS, p.30], it is enough to consider $\gamma\in A_{0}$, in which case
$\gamma=\left(\begin{array}[]{cc}\lambda&0\\ 0&\lambda^{-1}\end{array}\right)$
and then
$D_{G}(\gamma)=(\lambda+\lambda^{-1})^{2}-4=(\lambda-\lambda^{-1})^{2}.$
We also have
$D^{G}(\gamma)$ $=$ $|\Pi_{\alpha}(\alpha(\gamma)-1)|^{1/2}$
$=$ $|(\lambda^{2}-1)(\lambda^{-2}-1)|^{1/2}$
$=$ $|\lambda(\lambda-\lambda^{-1})(\lambda^{-1}-\lambda)\lambda^{-1}|^{1/2}$
$=$ $|\lambda-\lambda^{-1}|.$
It follows that
(2.12) $|D_{G}(\gamma)|^{1/2}=D^{G}(\gamma).$
The term $|D_{G}(\gamma)|^{1/2}$ occurs repeatedly in the denominator of the character formulas in [ADSS]. Thanks to equation (2.12), we shall write this denominator as $D^{G}(\gamma)$.
If
$\gamma=\left(\begin{matrix}a&b\cr b\varepsilon&a\end{matrix}\right)$
then we have
$D_{G}(\gamma)$ $=$ $4\displaystyle 4a^{2}-$ 4
$=$ $4\varepsilon b^{2},$
so that
$D^{G}(\gamma)=|b|.$
2.0.3. The function $f$
We set
(2.13) $f(\gamma):=\frac{\mathrm{sgn}_{\varepsilon}(\mathrm{Im}_{\varepsilon}(\gamma))}{|D_{G}(\gamma)|^{1/2}}.$
Note that
$\gamma\quad\textrm{is regular}\iff b\neq 0.$
In that case, we have
$f(\gamma)$ $=$ $\frac{\mathrm{sgn}_{\varepsilon}(b)}{|b|}$
$=$ $(-1)^{\mathrm{v}(b)}q^{\mathrm{v}(b)}$
$=$ $(-q)^{\mathrm{v}(b)}.$
It is worth noting that $f$ is invariant under the Galois group $\mathrm{Gal}(F(\sqrt{\varepsilon})/F)$. If $c=a+b\sqrt{\varepsilon}$ then we have
(2.14) $f(\overline{c})=f(a-b\sqrt{\varepsilon})=(-q)^{\mathrm{v}(-b)}=(-q)^{\mathrm{v}(b)}=f(c).$
Note that $D_{G}(\mathbf{g}\cdot\gamma)=D_{G}(\gamma)$ and so we have
$f(\mathbf{g}\cdot\gamma)=f(\gamma).$
2.0.4. Filtration
We refer to §5.1 in [ADSS]. Let $r=1$. Then we have
$\mathfrak{gl}_{2}(F)_{x_{L},1}=\left(\begin{array}[]{cc}\mathfrak{p}_{F}&\mathfrak{p}_{F}\\ \mathfrak{p}_{F}&\mathfrak{p}_{F}\end{array}\right)$
and
$G_{x_{L},1}$ $=$ $\mathrm{SL}_{2}(F)\cap(1+\mathfrak{gl}_{2}(F)_{x_{L},1})$
$=$ $\mathrm{SL}_{2}(F)\cap\left(\begin{array}[]{cc}1+\mathfrak{p}_{F}&\mathfrak{p}_{F}\\ \mathfrak{p}_{F}&1+\mathfrak{p}_{F}\end{array}\right)$
and
$G_{0+}\cap T^{\varepsilon}$ $=$ $G_{x_{L},1}\cap T^{\varepsilon}$
$=$ $\left(\begin{array}[]{cc}1+\mathfrak{p}_{F}&\mathfrak{p}_{F}\\ \mathfrak{p}_{F}&1+\mathfrak{p}_{F}\end{array}\right)\cap T^{\varepsilon}$
$=$ $T_{1}^{\varepsilon}.$
This notation is consistent with the notation in §5.2 of [ADSS]. The filtration of the elliptic torus $T^{\varepsilon}$ is associated to a unique point in the building of $G$ over $F$, namely $x_{L}$. In §5.2 we have
$T^{\varepsilon}_{r}=T^{\varepsilon}\cap G_{x_{L},r}$
and our equation arises when $r=1$.
Note that, again referring to §5.1 in [ADSS],
(2.17) $\gamma\in T_{1}^{\varepsilon}\implies\gamma\in G_{x_{L},1}\implies\mathbf{g}\cdot\gamma\in G_{x_{R},1}\subset G_{0+}$
We shall need (2.17) in the proof of Theorem 5.1.
The compact open subgroup $T^{\varepsilon}_{1}$ is the first compact open subgroup in the Moy-Prasad filtration of $T^{\varepsilon}$, see [Au1, 4.3.7.9].
The following Lemma is proved in [ADSS, p.65].
Lemma 2.18.
Let $\gamma$ be a regular semisimple element in the elliptic torus $T^{\varepsilon}$ which is far from the identity. Then we have $f(\gamma)=1$.
3. Depth-zero supercuspidal representations of $\mathrm{SL}_{2}(F)$
Let $I$ denote a line-segment (chamber) in the reduced building (tree) of $\mathrm{SL}_{2}(F)$. The ends of the interval $I$ will be denoted $x_{0}$, $x_{1}$ The stabilizer of $x_{i}$ will be denoted $G_{x_{i}}$, for $i=0,1$. We have $G_{x_{0}}=\mathrm{SL}_{2}(\mathfrak{o}_{F})$, and $G_{x_{1}}$ is the conjugate of $G_{x_{0}}$ under the element $\mathbf{g}$ of $\mathrm{GL}_{2}(F)$ defined in (2.5). The reductive quotients of $G_{x_{0}}$ and $G_{x_{1}}$ are both isomorphic to $\mathrm{SL}_{2}(\mathfrak{f})$, and $\{G_{x_{0}},G_{x_{1}}\}$ is a set of representatives of the maximal parahoric subgroups of $\mathrm{SL}_{2}(F)$.
We denote by $\mathbb{T}^{\epsilon}$ the unique maximal elliptic $\mathfrak{f}$-torus in $\mathrm{SL}_{2/\mathfrak{f}}$. Similarly to (2.1), we have
(3.1) $\mathbb{T}^{\epsilon}(\mathfrak{f})=\left\{\left(\begin{matrix}a&b\\ \epsilon b&a\end{matrix}\right)\,:\,\text{$a,b\in\mathfrak{f}$ and $a^{2}-\epsilon b^{2}=1$}\right\}.$
Let $\psi$ be a character of $\mathbb{T}^{\epsilon}(\mathfrak{f})$ such that $\psi\neq\psi^{-1}$. The Deligne-Lusztig virtual character $R_{\mathbb{T}^{\epsilon},\psi}^{\mathrm{SL}_{2}}$ is irreducible, and we denote by $|R_{\mathbb{T}^{\epsilon},\psi}|$ its inflation to $G_{x_{0}}$ and $G_{x_{1}}$. We put
(3.2) $\pi^{+}=\pi(\mathbb{T}^{\epsilon},\psi):=\mathrm{Ind}_{G_{x_{0},0}}^{G}|R^{G}_{\mathbb{T}^{\epsilon},\psi}|\quad\text{and}\quad\pi^{-}:=\mathrm{Ind}_{G_{x_{1},0}}^{G}|R^{G}_{\mathbb{T}^{\epsilon},\psi}|.$
The two compactly induced representations $\mathrm{c\!\!-\!\!Ind}_{G_{x_{i}}}^{G}|R_{\mathbb{T}^{\epsilon},\psi}|$, $i=0,1$, are depth-zero regular supercuspidal irreducible representations of $G$ in the terminology of [Ka2].
Let $\psi_{0}$ be the quadratic character of $\mathbb{T}^{\epsilon}(\mathfrak{f})$. The Deligne-Lusztig virtual character $R_{\mathbb{T}^{\epsilon},\psi_{0}}^{\mathrm{SL}_{2}}$ decomposes as
(3.3) $R_{\mathbb{T}^{\epsilon},\psi_{0}}^{\mathrm{SL}_{2}}=-R_{\mathbb{T}^{\epsilon},\psi_{0}}^{+}-R_{\mathbb{T}^{\epsilon},\psi_{0}}^{-},$
where $R_{\mathbb{T}^{\epsilon},\psi_{0}}^{\pm}$ are irreducible cuspidal representations of $\mathrm{SL}_{2}(\mathfrak{f})$ of dimension $\frac{q-1}{2}$ (see [DM, Table 12.1]). We choose the signs as in [ADSS, (9.3)].
We denote by $|R_{\mathbb{T}^{\epsilon},\psi_{0}}^{\pm}|$ the inflations of $R_{\mathbb{T}^{\epsilon},\psi_{0}}^{\pm}$ to $G_{x_{0}}$ and $G_{x_{1}}$. The four compactly induced representations $\mathrm{c\!\!-\!\!Ind}_{G_{x_{i}}}^{G}|R_{\mathbb{T}^{\epsilon},\psi_{0}}^{\pm}|$, $i=0,1$, are depth-zero supercuspidal irreducible representations of $G$, they are called “exceptional” in [ADSS], and are non-singular non-regular in the terminology of [Ka3] (see also [Au2]). We write
(3.4) $\pi^{+}:=\mathrm{c\!\!-\!\!Ind}_{G_{x_{0}}}^{G}|R_{\mathbb{T}^{\epsilon},\psi_{0}}^{+}|\quad\text{and}\quad\pi^{-}:=\mathrm{c\!\!-\!\!Ind}_{G_{x_{1}}}^{G}|R_{\mathbb{T}^{\epsilon},\psi_{0}}^{-}|.$
We define
$(\mathbf{g}\cdot\pi^{+})(x):=\pi^{+}(\mathbf{g}^{-1}x\mathbf{g})\quad\text{and}\quad\mathbf{g}\cdot x:=\mathbf{g}x\mathbf{g}^{-1},$
so that
$(\mathbf{g}\cdot\pi^{+})(x)=\pi^{+}(\mathbf{g}^{-1}\cdot x).$
We set
(3.5) $\pi_{1}=\pi^{+},\quad\pi_{2}=\pi^{-},\quad\pi_{3}=\mathbf{g}\cdot\pi^{+},\quad\pi_{4}=\mathbf{g}\cdot\pi^{-}.$
4. Virtual characters far from the identity: the non-regular parameter
Let $\phi$ be the Langlands parameter (7.1). We consider the associated supercuspidal $L$-packet.
$\Pi_{\phi}(\mathrm{SL}_{2}(F))=\{\pi_{1},\pi_{2},\pi_{3}.\pi_{4}\},$
and let $\Theta_{j}:=\Theta_{\pi_{j}}$ with $1\leq j\leq 4$.
Theorem 4.1.
Let $\gamma$ be a regular semisimple element in the elliptic torus $T^{\varepsilon}$ which is far from the identity. Then we have the character formulas
$\Theta_{1}(\gamma)$ $=$ $-\psi_{0}(\gamma)$
$\Theta_{2}(\gamma)$ $=$ $-\psi_{0}(\gamma)$
$\Theta_{3}(\gamma)$ $=$ 00 0
$\Theta_{4}(\gamma)$ $=$ $0$
Proof.
We refer to equations ()(*)( ∗ ) and ()(**)( ∗ ∗ ) on p.65 of [ADSS]. It follows immediately that
$\Theta_{\pi^{+}}(\gamma)=\Theta_{\pi^{-}}(\gamma)=-\psi_{0}(\gamma).$
We have, for all $\gamma\in T^{\varepsilon}$,
$\Theta_{\pi_{3}}(\gamma)$ $=$ $\Theta_{g\cdot\pi^{+}}(\gamma)$
$=$ $\Theta_{\pi^{+}}(g^{-1}\cdot\gamma)$
$=$ 00 0
since $\Theta_{\pi^{+}}$ vanishes on $T^{\varepsilon,\varpi}$ by [ADSS, Theorem 5.1].
Similarly, we have, for all $\gamma\in T^{\varepsilon}$,
$\Theta_{\pi_{4}}(\gamma)$ $=$ $\Theta_{g\cdot\pi^{-}}(\gamma)$
$=$ $\Theta_{\pi^{-}}(g^{-1}\cdot\gamma)$
$=$ 00 0
since $\Theta_{\pi^{-}}$ vanishes on $T^{\varepsilon,\varpi}$ by [ADSS, Theorem 5.1]. ∎
Consider the $L$-packet of $\mathrm{SL}_{2}(F)$ given by
(4.2) $\Pi_{\phi}(\mathrm{SL}_{2}(F))=\{\pi_{1},\pi_{2},\pi_{3},\pi_{4}\}$
where the representations $\pi_{i}$ are depth-zero supercuspidal irreducible representations of $G$ (see (3.5)). We suppose that $\pi_{1}$ is $\mathfrak{w}$-generic.
Let $\Theta_{1}$, $\Theta_{2}$, $\Theta_{3}$, $\Theta_{4}$ be the distribution characters of $\pi_{1}$, $\pi_{2}$, $\pi_{3}$, $\pi_{4}$.
Given $A\in\mathrm{GL}_{2}(\mathbb{C})$ let $A_{*}$ denote the image of $A$ in $\mathrm{PGL}_{2}(\mathbb{C})$. Let
(4.3) $s_{1}=\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right)_{*}\quad s_{2}=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right)_{*}\quad s_{3}=\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)_{*}.$
We have
$S_{\phi}=\{1,s_{1},s_{2},s_{1}s_{2}\}.$
We will enumerate the characters of $S_{\phi}$ in the following way:
$1$ $s_{1}$ $s_{2}$ $s_{1}s_{2}$
$\rho_{1}$ $1$ $1$ $1$ $1$
$\rho_{2}$ $1$ $1$ $-1$ $-1$
$\rho_{3}$ $1$ $-1$ $1$ $-1$
$\rho_{4}$ $1$ $-1$ $-1$ $1$
Given
(4.4) $\iota\colon\Pi_{\phi}\to\mathrm{Irr}(S_{\phi})$
we attach to $(\phi,s)$ the virtual character
(4.5) $\Theta_{\phi,s}:=\sum_{\pi\in\Pi_{\phi}(\mathrm{SL}_{2}(F))}\iota(\pi)(s)\cdot\Theta_{\pi}.$
The first constraint is that $\iota$ must assign the trivial character $\rho_{1}$ to the $\mathfrak{w}$-generic representation $\pi_{1}$. The second important constraint is the endoscopic character identity. These constraints determine the map $\iota$ uniquely. The defining equation is
(4.6) $\iota(\pi_{j}):=\rho_{j}\quad\text{ with $j=1,2,3,4$.}$
In that case, we have
(4.7) $\Theta_{\phi,s}=\begin{cases}\Theta_{1}+\Theta_{2}+\Theta_{3}+\Theta_{4}&\text{if $s=1$}\\ \Theta_{1}+\Theta_{2}-\Theta_{3}-\Theta_{4}&\text{if $s=s_{1}$}\\ \Theta_{1}-\Theta_{2}+\Theta_{3}-\Theta_{4}.&\text{if $s=s_{2}$}\\ \Theta_{1}-\Theta_{2}-\Theta_{3}+\Theta_{4}&\text{if $s=s_{3}$}\end{cases}.$
Theorem 4.8.
Let $\gamma$ be a regular semisimple element in the elliptic torus $T^{\varepsilon}$ which is far from the identity. Then we have
$\Theta_{\phi,1\hphantom{s}}(\gamma)$ $=$ $-2\psi_{0}(\gamma)$
$\Theta_{\phi,s_{1}}(\gamma)$ $=$ $-2\psi_{0}(\gamma)$
$\Theta_{\phi,s_{2}}(\gamma)$ $=$ 00 0
$\Theta_{\phi,s_{3}}(\gamma)$ $=$ $0$
Proof.
This follows readily from Theorem 4.1. ∎
5. Virtual characters near the identity: the non-regular parameter
The sign character $\mathrm{sgn}_{\varepsilon}\colon F^{\times}\to\{\pm 1\}$ is determined by the short exact sequence
$1\to{\mathrm{N}}_{E/F}E^{\times}\to F^{\times}\to\mathrm{Gal}(E/F)\to 1$
The notation $\mathrm{Im}_{\varepsilon}$ has already been defined in §3.1. We recall the definition of $f(\gamma)$ from (2.13). Throughout this section, $\gamma$ is near the identity.
We will use the Cayley transform defined on an open subset of $\mathfrak{g}$ by
$c\colon X\mapsto\frac{1+X/2}{1-X/2}.$
We recall from [ADSS, Lemma 5.5(h)] that $|D_{\mathfrak{g}}(Y)|=|D_{G}(c(Y))|$ when $Y\in\mathfrak{g}_{0+}$.
Theorem 5.1.
Let $\gamma$ be a regular semisimple element in the elliptic torus $T^{\varepsilon}$ which is near the identity. Then we have
$\Theta_{\phi,1}(\gamma)$ $=$ $-2$
$\Theta_{\phi,s_{1}}(\gamma)$ $=$ $-2f(\gamma).$
Proof.
Since $\gamma\in G^{\mathrm{rss}}\cap G_{0+}$, according to Proposition 13.14 in [ADSS], we have
(5.2) $\Theta_{1}(\gamma)+\Theta_{2}(\gamma)=\hat{\mu}^{G}_{X_{1}^{\varepsilon,1}}(c^{-1}(\gamma)).$
We will use definition ($\ddagger$) in [ADSS, p.65–66] and first take $\theta^{\prime}=\varepsilon$, $\eta=1$. Recall the definition of $T^{\varepsilon}$ from (2.2). With $Y$ in the Lie algebra of $T^{\varepsilon}$, by using (2.12), we then have
$[A;B_{1},B_{\varepsilon},B_{\varpi},B_{\varepsilon\varpi}]_{\varepsilon,0}(Y):=|\varepsilon|^{1/2}A+q^{-1}\frac{1}{D^{\mathfrak{g}}(Y)}B_{\varepsilon}(\mathrm{Im}_{\varepsilon}(Y)).$
In particular, we have
$[-1;1,B_{\varepsilon},0,0]_{\varepsilon,0}(Y)=-1+q^{-1}\frac{1}{D^{\mathfrak{g}}(Y)}B_{\varepsilon}(\mathrm{Im}_{\varepsilon}(Y)).$
According to ()(*)( ∗ ) on p.66 of [ADSS], we have 111we made a (surely necessary) correction here
(5.3) $B_{\varepsilon}=q\cdot H(\Lambda^{\prime},F(\sqrt{\varepsilon}))\mathrm{sgn}_{\varepsilon}\\ =-q\cdot\mathrm{sgn}_{\varepsilon}$
since $H(\Lambda^{\prime},F(\sqrt{\varepsilon}))=-1$, by [ADSS, Lemma 4.2].
Taking $Y=c^{-1}(\gamma)$, we obtain
$\hat{\mu}^{G}_{X_{1}^{\varepsilon,1}}(c^{-1}(\gamma))$ $=$ $[-1;1,B_{\varepsilon},0,0]_{\varepsilon,0}$
$=$ $-1-\frac{\mathrm{sgn}_{\varepsilon}(\mathrm{Im}_{\varepsilon}(\gamma))}{D^{G}(\gamma)}$
$=$ $-1-f(\gamma)$
by (2.13). Hence, (5.2) gives
(5.4) $\Theta_{1}(\gamma)+\Theta_{2}(\gamma)=-1-f(\gamma).$
We now take $Y=c^{-1}(\mathbf{g}\cdot\gamma)$. Since $Y$ is in the Lie algebra of $T^{\varepsilon,\varpi}$ (by (2.6)), we take $\theta^{\prime}=\varepsilon$, $\eta=\varpi$ in the formula in [ADSS, 3rd line of p.66] and $B_{\varepsilon}$ as in (5.3). We obtain
(5.5) $[-1;1,B_{\varepsilon},0,0]_{\varepsilon,0}(Y)$ $=$ $-1+q^{-1}\frac{1}{D^{\mathfrak{g}}(Y)}B_{\varepsilon}(\varpi^{-1}\mathrm{Im}_{\varepsilon}(Y))$
(5.6) $=$ $-1+f(\gamma).$
According to ()(*)( ∗ ) in [ADSS, §15.2], we have
(5.7) $\hat{\mu}^{G}_{X_{1}^{\varepsilon,1}}(c^{-1}(\mathbf{g}\cdot\gamma))=[-1;1,B_{\varepsilon},0,0]_{\varepsilon,0}(Y).$
It follows from (5.5) and (5.7) that
(5.8) $\hat{\mu}^{G}_{X_{1}^{\varepsilon,1}}(c^{-1}(\mathbf{g}\cdot\gamma))=-1+f(\gamma).$
We note that $\mathbf{g}\cdot\gamma\in G^{\mathrm{rss}}\cap G_{0+}$ by (2.17). Now, using [ADSS, Theorem 13.14] and (5.8), we obtain
(5.9) $\Theta_{3}(\gamma)+\Theta_{4}(\gamma)$ $=$ $\Theta_{1}(\mathbf{g}\cdot\gamma)+\Theta_{2}(\mathbf{g}\cdot\gamma)$
(5.10) $=$ $\hat{\mu}^{G}_{X_{1}^{\varepsilon,1}}(c^{-1}(\mathbf{g}\cdot\gamma))$
(5.11) $=$ $-1+f(\gamma).$
Therefore, using (5.4) and (5.9), we have
$\Theta_{\phi,s_{1}}$ $=$ $\Theta_{1}+\Theta_{2}-\Theta_{3}-\Theta_{4}$
$=$ $-1-f(\gamma)-(-1+f(\gamma))$
$=$ $-2f(\gamma)$
and
$\Theta_{\phi,1}$ $=$ $\Theta_{1}+\Theta_{2}+\Theta_{3}+\Theta_{4}$
$=$ $-1-f(\gamma)-1+f(\gamma)$
$=$ $-2$
as required. ∎
Problems with Theorem 15.2 in [ADSS]. There are two inconsistencies. The first inconsistency is between Proposition 13.14 and Theorem 15.2 in [ADSS]. We consider the sum $\Theta_{1}(\gamma)+\Theta_{2}(\gamma)=\Theta_{\pi^{+}}(\gamma)+\Theta_{\pi^{-}}(\gamma)$ by (3.5).
According to Theorem 15.2 in [ADSS] with $\theta^{\prime}=\varepsilon$ and $\eta=1$ we have
(5.12) $\Theta_{1}(\gamma)+\Theta_{2}(\gamma)=-1$
So there is an obvious clash between (5.4) and (5.12).
The second inconsistency is between the following consequence of Theorem 15.2 in [ADSS]) and the endoscopic character identity.
Consequence of Theorem 15.2. If $\gamma\in T^{\varepsilon}_{1}$ then we have
$\Theta_{1}(\gamma)$ $=$ $\frac{-f(\gamma)-1}{2}$
$\Theta_{2}(\gamma)$ $=$ $\frac{f(\gamma)-1}{2}$
$\Theta_{3}(\gamma)$ $=$ $\frac{f(\gamma)-1}{2}$
$\Theta_{4}(\gamma)$ $=$ $\frac{-f(\gamma)-1}{2}.$
and
$\Theta_{\phi,1}(\gamma)$ $=$ $-2$
$\Theta_{\phi,s_{1}}(\gamma)$ $=$ 00 0
$\Theta_{\phi,s_{2}}(\gamma)$ $=$ 00 0
$\Theta_{\phi,s_{3}}(\gamma)$ $=$ $-2f(\gamma).$
This clashes with the endoscopic character identity.
Proof.
The formula for $\Theta_{1}(\gamma)$ follows from [ADSS, Theorem 15.2] with, in their notation, $\theta^{\prime}=\varepsilon$ and $\eta=1$. We have used the fact that, in their notation, we have $H(\Lambda^{\prime},F(\sqrt{\varepsilon}))=-1$. This depends on [ADSS, Lemma 4.2] and the fact that $\Lambda^{\prime}$ is an additive character of depth zero.
The formula for $\Theta_{2}(\gamma)$ follows from loc. cit., again with $\theta^{\prime}=\varepsilon$ and $\eta=1$.
Moving on to the character $\Theta_{3}$, we have $\Theta_{3}(\gamma)=\Theta_{1}(\mathbf{g}\cdot\gamma)$. Since $\gamma\in T^{\varepsilon}$, we have $\mathbf{g}\cdot\gamma\in T^{\varepsilon,\varpi}$, by (2.6). Thus we get
(5.13) $\Theta_{3}(\gamma)=\frac{1}{2}\left\{\frac{-\mathrm{sgn}_{\varepsilon}(\varpi^{-1}\mathrm{Im}_{\varepsilon}(\mathbf{g}\cdot\gamma))}{D^{G}(\gamma)}-1\right\}$
using Theorem 15.2 in [ADSS] with $\theta^{\prime}=\varepsilon$ and $\eta=\varpi$. We now apply (2.7) and infer that
$\Theta_{3}(\gamma)=(f(\gamma)-1)/2.$
A parallel argument leads us to
$\Theta_{4}(\gamma)=(-f(\gamma)-1)/2.$
Recall (4.7). Then the last part of the theorem now follows rather readily. For example:
$\Theta_{\phi,s_{3}}$ $=$ $\Theta_{1}-\Theta_{2}-\Theta_{3}+\Theta_{4}$
$=$ $-f(\gamma)-f(\gamma)$
$=$ $-2f(\gamma)$
6. Virtual characters: regular parameters
Theorem 6.1.
Let $s$ be the non-trivial element in $S_{\phi}$. Let $\gamma$ be a regular semisimple element in the elliptic torus $T^{\varepsilon}$ which is far from the identity. Then we have
$\Theta_{\phi,s}(\gamma)$ $=$ $-\psi(\gamma^{H})-\psi(1/\gamma^{H}).$
Proof.
The equation
$\Theta_{\pi}(\gamma)=-\psi(\gamma^{H})-\psi(1/\gamma^{H})$
is a re-statement of Theorem 14.14 in [ADSS]. We have simplified this statement by using Lemma 2.18 and [ADSS, Lemma 4.2].
According to [ADSS, Theorem 14.14] we have
$\Theta_{\pi^{-}}(\gamma)=0.$
In that case, we have
$\Theta_{\phi,s}(\gamma)$ $=$ $\sum_{\pi\in\Pi_{\phi}(\mathrm{SL}_{2}(F))}\mathrm{trace}(\iota(\pi))(s)\cdot\Theta_{\pi}$
$=$ $\rho_{0}(s)\Theta_{\pi}(\gamma)+\rho_{1}(s)\Theta_{\pi^{-}}(\gamma)$
$=$ $\Theta_{\pi}(\gamma)$
$=$ $-\psi(\gamma^{H})-\psi(1/\gamma^{H}).$
Theorem 6.2.
Let $s$ be the non-trivial element in $S_{\phi}$. Let $\gamma$ be a regular semisimple element in the elliptic torus $T^{\varepsilon}$ which is near the identity. Then we have
$\Theta_{\phi,s}(\gamma)$ $=$ $-2f(\gamma)$
Proof.
Let the depth-zero supercuspidal representation $\pi$ be defined as
$\pi:=\pi(T^{\varepsilon},\psi)$
as in [ADSS, Definition 9.6] (which is the representation $\pi^{+}$ defined in (3.2)). In the terminology of [ADSS, Definition 14.1], we are in the unramified case. Recall that
$\pi^{-}=\mathbf{g}\cdot\pi^{+}.$
The $L$-packet is then
$\Pi_{\phi}=\{\pi^{+},\pi^{-}\}.$
According to [ADSS, Theorem 14.20] we have
$\gamma\in T_{1}^{\varepsilon}\implies\Theta_{\pi}(\gamma)=c_{0}(\pi)-f(\gamma).$
By [ADSS, Definition 14.17], the constant term $c_{0}(\pi)$ is
$c_{0}(\pi)=-1$
and so
$\Theta_{\pi}(\gamma)=-1-f(\gamma).$
Note that $\gamma\in T^{\varepsilon}\implies\mathbf{g}\cdot\gamma\in T^{\varepsilon,\varpi}$. With $\gamma\in T_{1}^{\varepsilon}$ we have
$\Theta_{\pi}(\mathbf{g}\cdot\gamma)$ $=$ $c_{0}(\pi)-\frac{\mathrm{sgn}_{\varepsilon}(\varpi^{-1}\mathrm{Im}_{\varepsilon}(\mathbf{g}\cdot\gamma))}{D^{G}(\mathbf{g}\cdot\gamma)}$
by Theorem 14.20 in [ADSS]. Using 2.7, we obtain
$\Theta_{\pi}(\mathbf{g}\cdot\gamma)=c_{0}(\pi)+f(\gamma)$
With $\gamma\in T^{\varepsilon}$ we also have
$\Theta_{\pi^{-}}(\gamma)$ $=$ $\Theta_{\pi}(\mathbf{g}\cdot\gamma)$
and so
$\Theta_{\pi^{-}}(\gamma)=-1+f(\gamma).$
We recall that $S_{\phi}=\{1,s\}=\mathbb{Z}/2\mathbb{Z}$. Denote the trivial character of $S_{\phi}$ by $\rho_{0}$ and the non-trivial character by $\rho_{1}$. Assume that $\pi$ is $\mathfrak{w}$-generic so that the parametrization of the $L$-packet is
$\iota\colon\Pi_{\phi}\to\mathrm{Irr}(S_{\phi}),\quad\quad\pi\mapsto\rho_{0},\quad\pi^{-}\mapsto\rho_{1}.$
The virtual character is given by
$\Theta_{\phi,s}(\gamma)$ $=$ $\rho_{0}(s)\Theta_{\pi}(\gamma)+\rho_{1}(s)\Theta_{\pi^{-}}(\gamma)$
$=$ $\Theta_{\pi}(\gamma)-\Theta_{\pi^{-}}(\gamma)$
$=$ $-1-f(\gamma)-(-1+f(\gamma))$
$=$ $-2f(\gamma).$
7. Endoscopy: the non-regular parameter
We recall that $G=\mathrm{SL}_{2}(F),\,\widehat{G}=\mathrm{PGL}_{2}(\mathbb{C})=\mathrm{PSL}_{2}(\mathbb{C})$. Let $E/F$, $E^{\prime}/F$ be distinct quadratic extensions, with $E/F$ unramified. Let $\sigma$ (resp. $\tau$) be the non-trivial element in $\mathrm{Gal}(E/F)$ (resp. $\mathrm{Gal}(E^{\prime}/F)$). Let $K/F$ denote the biquadratic extension comprising $E/F$ and $E^{\prime}/F$. Let $s,\mathbf{n}$ be a pair of commuting semisimple elements in $\widehat{G}$ of order $2$. The Langlands parameter
(7.1) $\phi\colon\mathbf{W}_{F}\to\mathrm{PGL}_{2}(\mathbb{C})$
is now defined by the condition that $\phi$ factors through $\mathrm{Gal}(K/F)$ and satisfies the equations
(7.2) $\phi(\sigma)$ $=$ $\mathbf{n},\quad\phi(\tau)=s.$
We now proceed to construct the endoscopic datum $\mathfrak{e}$. First, we need the quadratic character of the unramified elliptic torus.
Definition 7.3.
[ADSS, p.27]. Let $E^{1}$ denote the group of norm $1$ elements in $F(\sqrt{\varepsilon})$. For $\lambda\in E^{1}$ ,
$\psi_{0}(\lambda):=\begin{cases}\mathrm{sgn}_{\varpi}(\lambda+\lambda^{-1}+2)&\text{if $\lambda\neq-1$,}\\ -\mathrm{sgn}_{\varpi}(-1)&\text{if $\lambda=-1$.}\end{cases}$
Then $\psi_{0}$ is the unique quadratic character of $E^{1}$. The quadratic character $\psi_{0}$ factors through $T^{\varepsilon}/T^{\varepsilon}_{1}$ and so $\psi_{0}$ has depth zero [ADSS, §9.2]. If $\gamma$ is near the identity, then $\gamma\in T^{\varepsilon}_{1}$ and so $\psi_{0}(\gamma)=1$.
Let $s=s_{1}$, $\mathbf{n}=s_{2}$. Consider the identity component of the centralizer of $s$ in $\widehat{G}$. This is a complex torus of dimension $1$. Following [Ka5, §4.8] we will set $\widehat{H}=\mathrm{Cent}(s,\widehat{G})^{0}$ and consider $\mathcal{H}=\widehat{H}\cdot\phi(\mathbf{W}_{F})$ and the inclusion
$\xi:\mathcal{H}\to\widehat{G}$
Note that
$\widehat{H}\cdot\phi(\mathbf{W}_{F})=\widehat{H}\cdot\phi(\sigma)$
since $\phi(\tau)\in\widehat{H}$. Now $\phi(\sigma)$ acts on $\widehat{H}$ as inversion. For, let
$h=\left(\begin{array}[]{cc}1&0\\ 0&z\end{array}\right)_{*}\in\widehat{H}$
then we have
$\phi(\sigma)\cdot h$ $=$ $\mathbf{n}\left(\begin{array}[]{cc}1&0\\ 0&z\end{array}\right)_{*}\mathbf{n}^{-1}$
$=$ $\left(\begin{array}[]{cc}1&0\\ 0&1/z\end{array}\right)_{*}$
This determines an action of $\Gamma_{E/F}$ by inversion on $\widehat{H}$. Let $H$ be the quasi-split group defined over $F$ that is dual to $\widehat{H}$ and whose rational structure is determined by $\mathrm{Gal}(E/F)\to\mathrm{Out}(\widehat{H})=\mathrm{Out}(H)$ as above. Now the group $E^{1}$ admits a canonical action of the Galois group $\mathrm{Gal}(E/F)$, for if $e\in E^{1}$ then $\sigma(e)=\overline{e}=e^{-1}$. So $E^{1}$ qualifies as the endoscopic group $H$. This creates the endoscopic datum
$\mathfrak{e}=(H,\mathcal{H},s,\xi).$
We have
[…]
In the notation of [FKS] we have $(S,\theta)=(H,\psi_{0})$ and $\mathcal{H}={}^{L}S$. The map $\mathbf{W}_{F}\to{}^{L}S$, denoted $\phi^{H}$, is the $L$-parameter of $\psi_{0}$ and so we have
$\Theta_{\phi^{H}}=\psi_{0}$
and, for the stable character,
(7.6) $S\Theta_{\phi^{H}}=\psi_{0}.$
We need to compute the absolute transfer factor $\Delta[\mathfrak{w},\mathfrak{e}]$.
Lemma 7.7.
Let $\gamma$ be a regular semisimple element in $T^{\varepsilon}$. Then we have
$\Delta[\mathfrak{w},\mathfrak{e}](\gamma^{H},\gamma)=-f(\gamma).$
Proof.
Underlying the map $\widehat{\eta}\colon\widehat{H}\to\widehat{G}$ we have two isomorphisms of $F$-tori:
$T^{\varepsilon}$$\to$ $H$
$\gamma=\left(\begin{array}[]{cc}a&b\\ \varepsilon b&a\end{array}\right)$ $\mapsto$ $\gamma^{H}=a+b\sqrt{\varepsilon}.$
and
$\gamma=\left(\begin{array}[]{cc}a&b\\ \varepsilon b&a\end{array}\right)$ $\mapsto$ $1/\gamma^{H}=a-b\sqrt{\varepsilon}.$
These isomorphisms are admissible in the sense of [Ka2, §3]. We therefore have two related pairs:
(7.10) $(\gamma^{H},\gamma)\quad\textrm{and}\quad(1/\gamma^{H},\gamma).$
To compute the absolute transfer factor, we will quote the Kaletha formula in [Ka4, Example 3.6.9]. We first need to convert notations as follows:
$\gamma=\gamma^{H}=c=a+b\eta=a+b\sqrt{\varepsilon},\quad\psi=\Lambda,\quad\lambda(E/F,\psi)=\epsilon(1/2,\mathbb{C},\psi)$
where $\mathbb{C}$ admits the following $\mathbf{W}_{F}$-module structure: $\mathbf{W}_{F}$ acts on $\mathbb{C}$ via $\mathrm{Gal}(E/F)$ and the non-trivial element in $\mathrm{Gal}(E/F)$ acts as $-1$. Then the Kaletha formula is
(7.11) $\epsilon(1/2,\mathbb{C},\psi)\cdot\kappa_{E/F}\left(\frac{c-\overline{c}}{2\eta}\right)|c-\overline{c}|_{F}.$
We recall that
$D^{G}(\gamma)=|c-\overline{c}|_{F}=|b|.$
In order to match the normalization of transfer factors in the recent article [AK], we switch the position of the Weyl discriminant $|c-\overline{c}|_{F}$ from numerator to denominator. The transfer factor defined in [AK], for which there is a clean formulation of the endoscopic character identity, is the transfer factor in [Ka1, p.163] divided by $D^{G}(\gamma)^{2}$. This leads to the equation
(7.12) $\Delta[\mathfrak{w},\mathfrak{e}](c,c)=\epsilon(1/2,\mathbb{C},\psi)\cdot\kappa_{E/F}\left(\frac{c-\overline{c}}{2\eta}\right)\cdot|c-\overline{c}|_{F}^{-1}.$
It follows that
$\Delta[\mathfrak{w},\mathfrak{e}](c,c)$ $=$ $\epsilon(1/2,\mathbb{C},\psi)\cdot\mathrm{sgn}_{\varepsilon}(b)\cdot|b|^{-1}.$
Local class field theory allows us to replace the $\mathbf{W}_{F}$-module $\mathbb{C}$ with the sign character $\mathrm{sgn}_{\varepsilon}$. Denote this character by $\chi$. We infer that
$\epsilon(1/2,\mathbb{C},\psi)=\epsilon(\chi,1/2,\psi)$
the Tate local constant of $\chi$ as in [BH, §23.4]. Note that “$\Lambda$ has depth zero” in [ADSS] is the same as “$\psi$ has level one” in [BH]. We now apply the Proposition in [BH, §23.5] and we obtain
$\epsilon(1/2,\mathbb{C},\psi)$ $=$ $\epsilon(\chi,1/2,\psi)$
$=$ $\chi(\varpi)^{-1}$
$=$ $-1.$
Putting all this together, we conclude that
(7.13) $\Delta[\mathfrak{w},\mathfrak{e}](\gamma^{H},\gamma)=-f(\gamma).$
According to (2.14), $f$ is invariant under the action of $\mathrm{Gal}(E/F)$, from which we infer that
$\Delta[\mathfrak{w},\mathfrak{e}](1/\gamma^{H},\gamma)=-f(\gamma).$
Theorem 7.14.
Let $\phi$ be the non-regular supercuspidal Langlands parameter for $\mathrm{SL}_{2}$. Then we have
$\sum_{\delta\in H(F)}\Delta[\mathfrak{w},\mathfrak{e}](\delta,\gamma)\,S\Theta_{\phi^{H}}(\delta)=-2f(\gamma)\psi_{0}(\gamma^{H}).$
Proof.
We have
[…]
$=$ $\Delta[\mathfrak{w},\mathfrak{e}](\gamma^{H},\gamma)\psi_{0}(\gamma^{H})$
$+\Delta([\mathfrak{w},\mathfrak{e}](1/\gamma^{H},\gamma))\psi_{0}(1/\gamma^{H})$
[…]
$=$ $-2f(\gamma)\psi_{0}(\gamma^{H})\quad\textrm{since}\quad\psi_{0}=\psi_{0}^{-1}.$
Remark 7.15.
In [ADSS] there is a dichotomy between the cases when $\gamma$ is far from the identity on the one hand, and near the identity on the other hand. This dichotomy is reflected in the two formulas for the term $\Delta[\mathfrak{w},\mathfrak{e}](\gamma,\gamma^{H})S\Theta_{\phi^{H}}(\gamma^{H})$ in the endoscopic character identity. We have
$\Delta[\mathfrak{w},\mathfrak{e}](\gamma,\gamma^{H})S\Theta_{\phi^{H}}(\gamma^{H})=-\psi_{0}(\gamma^{H})$
when $\gamma$ is far from the identity and
$\Delta[\mathfrak{w},\mathfrak{e}](\gamma,\gamma^{H})S\Theta_{\phi^{H}}(\gamma^{H})=-f(\gamma)$
when $\gamma$ is near the identity.
8. Endoscopy: regular parameters
We recall the elements $s_{1}$, $s_{2}$ and $s_{3}$ in $\mathrm{PGL}_{2}(\mathbb{C})$ defined in (4.3). Let $E/F$ be the unramified quadratic extension and let $\sigma$ be the generator of $\mathrm{Gal}(E/F)$. Let $\xi$ be a character of $E^{\times}$ such that
$\xi\neq\xi^{\sigma}\quad\textrm{and}\quad\psi=\xi|_{E^{1}}.$
Then $(E/F,\xi)$ is an admissible pair in the sense of [BH, §18]. Let
$\rho_{\xi}=\mathrm{Ind}_{E}^{F}\,\xi$
be the induced $\mathbf{W}_{F}$-module. This is a $2$-dimensional imprimitive $\mathbf{W}_{F}$-module. We describe this module explicitly.
We note that $\mathbf{W}_{F}/\mathbf{W}_{E}\simeq\mathrm{Gal}(E/F)$. Coset representatives are $\{1,\sigma\}$.
As a vector space, the induced module is $\mathbb{C}\oplus\mathbb{C}$ with module structure as follows:
  • $w\in\mathbf{W}_{E}$ acts on the first summand $\mathbb{C}$ as $\xi(w)$
  • $w\in\mathbf{W}_{E}$ acts on the second summand $C$ as $\xi^{\sigma}(w)$
  • $\sigma$ sends $(z_{1},z_{2})$ to $(z_{2},z_{1})$.
This $\mathbf{W}_{F}$-module determines a homomorphism
$\rho_{\xi}\colon\mathbf{W}_{F}$$\to$ $\mathrm{GL}_{2}(\mathbb{C})$
$w$ $\mapsto$ $\left(\begin{array}[]{cc}\xi(w)&0\\ 0&\xi^{\sigma}(w)\end{array}\right)\quad\textrm{if}\quad w\in\mathbf{W}_{E}$
$\sigma$ $\mapsto$ $\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right).$
Then $\phi$ is the composition
[…]
The centralizer of $\mathrm{im}\,\phi$ in $\mathrm{PGL}_{2}(\mathbb{C})$ is $\{1,s_{1}\}$ and so we have
$S_{\phi}=\mathbb{Z}/2\mathbb{Z}.$
We consider the representations $\pi:=\pi^{+}=\pi(\mathbb{T}^{\epsilon},\psi)$ and $\pi^{-}$ defined in (3.2).
Lemma 8.4.
We have
$\Pi_{\phi}(\mathrm{SL}_{2}(F))=\{\pi,\pi^{-}\}.$
Proof.
Given the admissible pair $(E/F,\xi)$ the construction of the irreducible supercuspidal representation $\pi_{\xi}$ is described in [BH, 19.2]. Then we have the Langlands correspondence for $\mathrm{GL}_{2}(F)$ in [BH, 34.4]:
$\rho_{\xi}\mapsto\pi_{\Delta_{\xi}\xi}$
We note that, since $E/F$ is an unramified quadratic extension, the auxiliary character $\Delta_{\xi}$ is unramified of order $2$ by the Definition in [BH, 34.4]. Also $\kappa_{E/F}$ is unramified of order $2$ by the Proposition in [BH, 34.3]. Therefore we have $\Delta_{\xi}=\kappa_{E/F}$ and $\Delta_{\xi}$ has no effect on $\rho_{\xi}$ by the Lemma in [BH, 34.1]. So the Langlands correspondence is given by
$\rho_{\xi}\mapsto\pi_{\xi}.$
Now we refer to the definition of $\pi^{+}$ and $\pi^{-}$ in (3.2). ∎
Let $\rho_{0}$ denote the trivial character of $S_{\phi}$ and let $\rho_{1}$ denote the non-trivial character of $S_{\phi}$. We will assume that $\pi^{+}$ is $\mathfrak{w}$-generic. Then we have a bijection
$\iota\colon\Pi_{\phi}(\mathrm{SL}_{2}(F))\to\mathrm{Irr}(S_{\phi}),\quad\quad\iota(\pi^{+})=\rho_{0},\quad\iota(\pi^{-})=\rho_{1}.$
Let $s=s_{1}$. Let $\widehat{H}$ denote the identity component of the centralizer of $s$ is $\mathrm{PGL}_{2}(\mathbb{C})$. Define
$\mathcal{H}:=\widehat{H}\cdot\phi(W_{F})$
and let $\xi:\mathcal{H}\to\mathrm{PGL}_{2}(\mathbb{C})$ denote inclusion. We recall that $E$ is the unramified quadratic extension of $F$. Noting that $\phi(\mathbf{W}_{E})\subset\widehat{H}$, we have
$\mathcal{H}=\widehat{H}\cdot\phi(\sigma)$
Since
$\phi(\sigma)=\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)$
this determines an action of $\Gamma_{E/F}$ by conjugation on $\widehat{H}$. Let $H$ be the quasi-split group defined over $F$ that is dual to $\widehat{H}$ and whose rational structure is determined by $\Gamma_{E/F}\to\mathrm{Out}(\widehat{H})=\mathrm{Out}(H)$. This uniquely determines $H$ as the group of norm $1$ elements in $F(\sqrt{\varepsilon})$. This creates the endoscopic datum
$\mathfrak{e}=(H,\mathcal{H},s,\xi).$
We note that this endoscopic datum is identical to the one constructed for the non-regular parameter. In the two cases (regular parameter, non-regular parameter) the endoscopic data are the same, but the “little parameters ” $\phi^{H}$ are different.
Theorem 8.5.
Let $\gamma$ be a regular semisimple element in $\mathrm{SL}_{2}(F)$. We have
$\sum_{\delta\in H(F)}\Delta[\mathfrak{w},\mathfrak{e}](\delta,\gamma)S\Theta_{\phi^{H}}(\delta)=-f(\gamma)(\psi(\gamma^{H})+\psi(1/\gamma^{H})).$
Proof.
We have
[…]
where the little map $\phi^{H}$ is the $L$-parameter of $\psi$ as in [Ka2, p.1148]. Since $\phi^{H}$ is the $L$-parameter of $\psi$ we have $\Theta_{\phi^{H}}=\psi$ and therefore
(8.6) $S\Theta_{\phi^{H}}=\psi.$
We have already fixed $\mathfrak{w}$. In that case, the conjecture in [KT, Conjecture 6.8(2)] for the parameter $\phi$ becomes
(8.7) $\Theta^{s}_{\phi}(\gamma)$ $=$ $\sum_{\delta\in H(F)}\Delta[\mathfrak{w},\mathfrak{e}](\delta,\gamma)S\Theta_{\phi^{H}}(\delta).$
We need to compute the absolute transfer factor $\Delta[\mathfrak{w},\mathfrak{e}](\delta,\gamma)$.
Let $\gamma$ be a regular semisimple element in $T^{\varepsilon}$. Admissible isomorphisms are defined in [Ka2, §5.1]. The elliptic torus $T^{\varepsilon}$ and the endoscopic group $H$ admit two admissible isomorphisms, namely
$T^{\varepsilon}$$\to$ $H$
$\gamma=\left(\begin{array}[]{cc}a&b\\ \varepsilon b&a\end{array}\right)$ $\mapsto$ $\gamma^{H}=a+b\sqrt{\varepsilon}$
and
$\gamma=\left(\begin{array}[]{cc}a&b\\ \varepsilon b&a\end{array}\right)$ $\mapsto$ $1/\gamma^{H}=a-b\sqrt{\varepsilon}$
We therefore have two related pairs:
(8.10) $(\gamma^{H},\gamma)\quad\textrm{and}\quad(1/\gamma^{H},\gamma).$
We recall that the Whittaker datum $\mathfrak{w}$ is fixed throughout this article. Now the constructions in this section fit into the framework of Example 3.6.9 in [Ka4]. In particular, the $L$-embedding ${}^{L}\eta$ sends the non-trivial element $\sigma\in\mathrm{Gal}(E/F)$ to the matrix
$\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right).$
It follows that the transfer factor is given by the formula (7), namely
(8.11) $\Delta[\mathfrak{w},\mathfrak{e}](\gamma^{H},\gamma)=\Delta[\mathfrak{w},\mathfrak{e}](1/\gamma^{H},\gamma)=-f(\gamma).$
Finally, we have
[…]
$=$ $\Delta[\mathfrak{w},\mathfrak{e}](\gamma^{H},\gamma)\psi(\gamma^{H})$
[…]
as required. ∎
We note that the endoscopic character identity
$\Theta_{\phi,s}(\gamma)=\sum_{\delta\in H(F)}\Delta[\mathfrak{w},\mathfrak{e}](\delta,\gamma)S\Theta_{\phi^{H}}(\delta)$
is essentially the character identity 2.5.2 in [Ka4].
9. Stability across the inner forms of $\mathrm{SL}_{2}$
Let $\phi$ denote the non-regular parameter of $\mathrm{SL}_{2}$. The finite group $S_{\phi}$ admits a pull-back from the adjoint group $\mathrm{PGL}_{2}(\mathbb{C})$ to its simply-connected cover $\mathrm{SL}_{2}(\mathbb{C})$. The pull-back group $\widetilde{S}_{\phi}$ is the quaternion group $Q$, which admits four one-dimensional representations $\rho_{1},\rho_{2},\rho_{3},\rho_{4}$ and one two-dimensional representation $\rho_{5}$. We have a bijective map
$\mathrm{Irr}(\widetilde{S}_{\phi})\simeq\{\pi_{1},\pi_{2},\pi_{3},\pi_{4},\pi_{5}\}.$
The $L$-packet $\Pi_{\phi}$ is therefore part of a compound packet, comprising five representations, where $\pi_{5}$ is a supercuspidal representation of the inner form $\mathrm{SL}_{1}(D)$. Let $\Theta_{5}$ denote the Harish-Chandra character of $\pi_{5}$.
Theorem 9.1.
The character of the supercuspidal representation $\pi_{5}$ is given by
$\Theta_{5}(\delta)=\psi_{0}(\gamma)$
with $\delta\in\mathrm{SL}_{1}(D)$ and $\gamma\in\mathrm{SL}_{2}(F)$ whenever $\gamma,\delta$ are related.
Proof.
We will consider $\mathrm{GL}_{2}(F)$. Now $\mathrm{GL}_{2}(F)$ admits one inner twist, namely $\mathrm{GL}_{1}(D)$ where the division algebra $D$ has the Brauer invariant
$\mathrm{inv}_{F}(D)=1/2\in\mathbb{Q}/\mathbb{Z}.$
Let $\pi_{0}$ be the depth-zero supercuspidal representation of $\mathrm{GL}_{2}(F)$ for which
$\pi_{0}|_{\mathrm{SL}_{2}(F)}=\pi_{1}\oplus\pi_{2}\oplus\pi_{3}\oplus\pi_{4}.$
Let $\mathrm{JL}$ denote the Jacquet-Langlands correspondence and let $\pi:=\mathrm{JL}^{-1}\pi_{0}$. The local Langlands correspondence for $\mathrm{GL}_{1}(D)$ is determined by the composition
[…]
The two elements in the set
$\{\pi,\mathrm{JL}(\pi\}$
therefore admit the same $L$-parameter $\phi$, and comprise a compound packet. The corresponding character identity is
$\Theta_{\pi}(\delta)=-\Theta_{\mathrm{JL}(\pi)}(\gamma)$
whenever $\gamma\longleftrightarrow\delta$.
We have, as a representation space for $Q\times\mathrm{SL}_{2}(F)$,
$V_{\pi(\phi)}=\bigoplus_{1\leq j\leq 4}\rho_{j}\otimes\pi_{j}$
and, as a representation space for $\mathrm{SL}_{2}(F)$,
$V_{\pi(\phi)}=\pi_{1}\oplus\pi_{2}\oplus\pi_{3}\oplus\pi_{4}.$
Now we have
$V_{\pi(\phi)}$ $=$ $\rho_{5}\otimes\pi_{5}$
as a representation space for the direct product $Q\times\mathrm{SL}_{1}(D)$, according to [HS, p.4–5]. Upon restriction to the subgroup $\{1\}\times\mathrm{SL}_{1}(D)$ we obtain
$V_{\pi(\phi)}=\rho_{5}(1)\otimes\pi_{5}$
as a representation space for $\mathrm{SL}_{1}(D)$. Now $\rho_{5}(1)$ is the identity operator on a complex vector space of dimension $2$, and so $\pi_{5}$ occurs with multiplicity $2$ in the restriction from $\mathrm{GL}_{1}(D)$ to $\mathrm{SL}_{1}(D)$. The character identity follows suit and we obtain
(9.2) $2\Theta_{5}(\delta)=-(\Theta_{1}+\Theta_{2}+\Theta_{3}+\Theta_{4})(\gamma)$
whenever $\gamma\longleftrightarrow\delta$ with $\gamma\in\mathrm{SL}_{2}(F)_{\mathrm{reg}},\delta\in\mathrm{SL}_{1}(D)_{\mathrm{reg}}$. Far from the identity, we have, by Theorem 4.8
(9.3) $\Theta_{1}+\Theta_{2}+\Theta_{3}+\Theta_{4}$ $=$ $-2\psi_{0}.$
Near the identity, we have, by Theorem 5.1
(9.4) $\Theta_{1}+\Theta_{2}+\Theta_{3}+\Theta_{4}$ $=$ $-2.$
These formulas are consistent because $\psi_{0}$ has depth zero.
By comparing (9.2) with (9.3) and (9.4) we conclude that the character of the supercuspidal representation $\pi_{5}$ is given by
$\Theta_{5}(\delta)=\psi_{0}(\gamma)$
with $\delta\in\mathrm{SL}_{1}(D)$ and $\gamma\in\mathrm{SL}_{2}(F)$ whenever $\gamma,\delta$ are related. ∎
The following corollary is a special case of Theorem 4.4.4(1) in [FKS].
Corollary 9.5.
The virtual character $S\Theta_{\phi,*}$ is stable across the two inner forms of $\mathrm{SL}_{2}$:
$S\Theta_{\phi,0}(\gamma)=S\Theta_{\phi,1}(\delta)$
for all stably conjugate strongly regular elements $\gamma\in\mathrm{SL}_{2}(F),\delta\in\mathrm{SL}_{1}(D)$.
Proof.
We have
$S\Theta_{\phi,0}$ $=$ $e(\mathrm{SL}_{2}(F))\cdot(\Theta_{1}+\Theta_{2}+\Theta_{3}+\Theta_{4})$
$=$ $\Theta_{1}+\Theta_{2}+\Theta_{3}+\Theta_{4}$
$=$ $-2\Theta_{5}$
$=$ $e(\mathrm{SL}_{1}(D))\cdot 2\Theta_{5}$
$=$ $S\Theta_{\phi,1}$
as required. ∎
[…]