### v3

# Comparison of the Sally-Shalika character formulas with the endoscopic character identities for $\mathrm{SL}_{2}$

Abstract.
We consider the depth-zero supercuspidal $L$-packets of $\mathrm{SL}_{2}(F)$ where $F$ is a non-archimedean local field of characteristic zero. We compare the explicit endoscopic character identities for $\mathrm{SL}_{2}(F)$ with the classical character formulas of Sally-Shalika. Our main result concerns the supercuspidal $L$-packet of size $4$. For this $L$-packet, we show how the norm $1$ groups $H_{1},H_{2},H_{3}$ in the three quadratic extensions of $F$ play a crucial role in the endoscopic character identities for $\mathrm{SL}_{2}(F)$.
Key words and phrases:
Endoscopy, characters, $p$-adic, $\mathrm{SL}_{2}$
1. Introduction
We consider the depth-zero supercuspidal $L$-packets of $\mathrm{SL}_{2}$ where $F$ is a non-archimedean local field of characteristic zero. We compare the explicit endoscopic character identities for $\mathrm{SL}_{2}$ with the classical character formulas of Sally-Shalika.
Our main result concerns the supercuspidal $L$-packet of size $4$. For this $L$-packet, we show how the norm $1$ groups $H_{1},H_{2},H_{3}$ in the three quadratic extensions of $F$ play a crucial role in the endoscopic character identities for $\mathrm{SL}_{2}(F)$.
The local field $F$ is a finite extension of $\mathbb{Q}_{p}$. We shall assume that $p>2$. Let $q=q_{F}$ denote the cardinality of the residue field of $F$. Let $\mathrm{v}(x)$ denote the valuation of $x\in F^{\times}$. Let $\mathfrak{o}_{F}$ denote the ring of integers of $F$, let $\varpi$ be a uniformiser for $F$, let $\mathfrak{p}_{F}=\varpi\mathfrak{o}_{F}$ be the maximal ideal of $\mathfrak{o}_{F}$, and let $\varepsilon$ be a fixed nonsquare element in $\mathfrak{o}_{F}^{\times}$. In that case, $\{1,\varpi,\varepsilon,\varepsilon\varpi\}$ is a set of representatives of $F^{\times}/(F^{\times})^{2}$.
Let $G=\mathrm{SL}_{2}(F)$, and let $\mathrm{Z}(G)$ denote the center of $G$. We shall focus on the elliptic torus
(1.1) $T^{\varepsilon}:=\left\{\left(\begin{matrix}a&b\cr b\varepsilon&a\end{matrix}\right)\;:\;a,b\in\mathfrak{o}_{F},a^{2}-\varepsilon b^{2}=1\right\}.$
This elliptic torus $T^{\varepsilon}$ is a representative of the single stable conjugacy class of unramified elliptic maximal $F$-tori in $\mathrm{SL}_{2}(F)$.
We will define $f$ as follows:
(1.2) $f\colon T^{\varepsilon}\to\mathbb{Z},\quad\quad\left(\begin{matrix}a&b\cr b\varepsilon&a\end{matrix}\right)\mapsto(-q)^{\mathrm{v}(b)}.$
We denote by $T^{\varepsilon}_{1}$ the pro-unipotent radical of $T^{\varepsilon}$, namely
(1.3) $T^{\varepsilon}_{1}:=\left\{\left(\begin{matrix}a&b\cr b\varepsilon&a\end{matrix}\right)\in T^{\varepsilon}\;:\;a\in 1+\mathfrak{p}_{F},\;b\in\mathfrak{p}_{F}\;\right\}.$
We will say, following [ADSS], that $\gamma$ is near the identity if $\gamma\in T_{1}^{\varepsilon}$, and far from the identity otherwise.
We have the unipotent subgroup
$U(F)=\left\{\left(\begin{matrix}1&x\cr 0&1\end{matrix}\right):x\in F\right\}\subset\mathrm{SL}_{2}(F)$
A Whittaker datum $\mathfrak{w}$ for $\mathrm{SL}_{2}(F)$ is a pair $(U,\theta)$ where $\theta:U(F)\to\mathbb{C}^{\times}$ is a non-trivial character. An irreducible smooth representation $(\pi,V)$ of $\mathrm{SL}_{2}(F)$ is called $\mathfrak{w}$-generic if the restriction of $\pi$ to $U(F)$ contains $\theta$. In this article, we choose and fix a Whittaker datum $\mathfrak{w}$ for $\mathrm{SL}_{2}(F)$.
If $G=\mathrm{SL}_{2}(F)$ then the Langlands dual group $\widehat{G}$ is $\mathrm{PGL}_{2}(\mathbb{C})$ and the $L$-group $^{L}G$ is $\mathrm{PGL}_{2}(\mathbb{C})\rtimes\mathbf{W}_{F}$. In the context of this article, $\phi$ will denote a Langlands parameter
$\phi:\mathbf{W}_{F}\to{}^{L}G.$
Let $S_{\phi}$ denote the centralizer in $\mathrm{PGL}_{2}(\mathbb{C})$ of the image of $\phi$, and $\mathrm{Irr}(S_{\phi})$ the set of isomorphism classes of irreducible representations of $S_{\phi}$. We have a bijective map
(1.4) $\iota_{\phi}\colon\Pi_{\phi}(\mathrm{SL}_{2}(F))\to\mathrm{Irr}(S_{\phi})$
that sends the $\mathfrak{w}$-generic representation of $\mathrm{SL}_{2}(F)$ to the trivial character of $S_{\phi}$.
Given $s\in S_{\phi}$, we attach to $(\phi,s)$ the virtual character
(1.5) $\Theta_{\phi,s}:=\sum_{\pi\in\Pi_{\phi}(\mathrm{SL}_{2}(F))}\mathrm{trace}(\iota_{\phi}(\pi))(s)\cdot\Theta_{\pi}.$
Let $\widehat{H}$ denote the identity component of the centralizer of $s$ in $\widehat{G}$. The extended endoscopic triple, as defined in [T, §6.2], is
$\mathfrak{e}(s):=(s,H,{}^{L}\eta).$
Then $\phi$ will factorize as follows:
$\begin{CD}\mathbf{W}_{F}@>{\phi^{H}}>{}>{}^{L}H@>{{}^{L}\eta}>{}>{}^{L}G.\end{CD}$
The map $\phi^{H}$ will be called the little parameter associated to $\phi$. This leads to the definition of the stable character. The stable character $\mathscr{S}\Theta_{\phi^{H}}$ is defined as
$\mathscr{S}\Theta_{\phi^{H}}:=\Theta_{\phi^{H},1}.$
Note that $H$ is an elliptic torus and the right-hand-side collapses to a single character of $H$.
Then the endoscopic character identity for $\mathrm{SL}_{2}$ is
(1.6) $\sum_{\delta\in H(F)}\Delta[\mathfrak{w},\mathfrak{e}(s)](\delta,\gamma)\mathscr{S}\Theta_{\phi^{H}}(\delta)=\Theta_{\phi,s}(\gamma)$
where $\Delta[\mathfrak{w},\mathfrak{e}(s)]$ is the Langlands-Shelstad transfer factor. In the context of quasi-split connected reductive groups, our reference for the endoscopic character identity is equation (6.4) in the article by Taibi [T]. From a purely notational point of view, we prefer to switch $\gamma$ and $\delta$ in Taibi’s equation (6.4).
The explicit endoscopic character identities, for the depth-zero supercuspidal $L$-packets of $\mathrm{SL}_{2}$, are stated in Theorems 1.8 and 1.9.
The characters of the discrete series for $\mathrm{SL}_{2}(F)$ were obtained by Sally-Shalika [SS]. In particular, for the depth-zero supercuspidal $L$-packets, the characters appear in Table 3 of [SS]. They are given as locally integrable functions on the maximal torus of $\mathrm{SL}_{2}(F)$ and on the elliptic tori in $\mathrm{SL}_{2}(F)$. Many years later, proofs were supplied by Adler, DeBacker, Sally and Spice in [ADSS]. For the most part, we find it more convenient to refer to this paper [ADSS].
Let $E$ denote the unramified quadratic extension $F(\sqrt{\varepsilon})$ and let $E^{1}$ denote the group of elements of norm $1$ in $E$. We have a canonical isomorphism of elliptic tori:
(1.7) $E^{1}\to T^{\varepsilon},\quad\quad a+b\sqrt{\varepsilon}\mapsto\left(\begin{matrix}a&b\cr b\varepsilon&a\end{matrix}\right)$
and we shall often identify $E^{1}$ with $T^{\varepsilon}$. A character $\psi$ of $E^{1}$ is quadratic if $\psi^{2}=1$ and non-quadratic if $\psi^{2}\neq 1$. The regular supercuspidal parameters are defined by Kaletha [Ka2, Def. 5.2.3]. The group $\mathrm{SL}_{2}(F)$ admits a unique non-regular parameter, which features in Theorem 1.9.
When we combine Theorems 6.1 and 8.5, we obtain the following result.
Theorem 1.8.
Let $\gamma$ be a regular, semisimple element in $T^{\varepsilon}$. Let $\Pi_{\phi}(\mathrm{SL}_{2}(F))$ be the depth-zero supercuspidal $L$-packet with regular $L$-parameter $\phi$. Let $s$ in (1.6) denote the non-trivial element in $S_{\phi}$. Then the endoscopic group $H$ is $E^{1}$. The little parameter $\phi^{H}$ is the Langlands parameter of a non-quadratic depth-zero character $\psi$ of $H$. We have
$\sum_{\delta\in H(F)}\Delta[\mathfrak{w},\mathfrak{e}(s)](\delta,\gamma)\mathscr{S}\Theta_{\phi^{H}}(\delta)$ $=$ $-f(\gamma)(\psi(\gamma)+\psi(1/\gamma))$
and
[…]
These equations are consistent because $f(\gamma)=1$ when $\gamma$ is far from the identity, and $\psi(\gamma)=1$ when $\gamma$ is near the identity.
The three quadratic extensions of $F$ are
$E_{1}=F(\sqrt{\varepsilon}),\quad\quad E_{2}=F(\sqrt{\varpi}),\quad\quad E_{3}=F(\sqrt{\varepsilon\varpi}).$
Let $H_{j}$ denote the group of norm $1$ elements in $E_{j}$ with $j=1,2,3$.
The left-hand-side of the endoscopic character identity depends on the element $s$ in $S_{\phi}$. In order to mark this dependence, we introduce the following definition:
$\mathscr{E}(s):=\sum_{\delta\in H}\Delta[\mathfrak{w},\mathfrak{e}(s)](\delta,-)\mathscr{S}\Theta_{\phi^{H}}(\delta)$
The value of $\mathscr{E}(s)$ at an element $\gamma$ will be denoted as follows:
$\mathscr{E}(s:\gamma):=\sum_{\delta\in H}\Delta[\mathfrak{w},\mathfrak{e}(s)](\delta,\gamma)\mathscr{S}\Theta_{\phi^{H}}(\delta)$
The field $F$ admits a unique biquadratic extension field $K/F$ and the non-regular parameter $\phi$ factors through the Galois group $\mathrm{Gal}(K/F)$, see §7. Let $s_{1},s_{2}$ and $s_{3}$ denote the images in $\mathrm{PGL}_{2}(\mathbb{C})$ of the matrices
$\left(\begin{matrix}1&0\\ 0&-1\end{matrix}\right),\quad\left(\begin{matrix}0&1\\ -1&0\end{matrix}\right),\quad\left(\begin{matrix}0&1\\ 1&0\end{matrix}\right).$
We attach to $s_{i}$ the extended endoscopic triple $\mathfrak{e}(s_{i})=(s_{i},H_{i},{}^{L}\eta_{i})$ where $^{L}\eta_{i}:{}^{L}H_{i}\to{}^{L}G$ extends the embedding $\widehat{H_{i}}\to\widehat{G}$.
When we combine Theorems 4.10, 5.1 and 7.10, we obtain the following result. An interesting feature of this result is that all three endoscopic groups $H_{1},H_{2},H_{3}$ are involved.
Theorem 1.9.
Let $\gamma$ be a regular, semisimple element in $T^{\varepsilon}$. Let $\Pi_{\phi}$ be the unique depth-zero supercuspidal $L$-packet with non-regular parameter $\phi$. Let $\psi_{0}$ denote the unique quadratic character of $T^{\varepsilon}$. We have
$\mathscr{E}(s_{1}:\gamma)$ $=$ $-2f(\gamma)\psi_{0}(\gamma)$
$\mathscr{E}(s_{2}:\gamma)$ $=$ $0$
$\mathscr{E}(s_{3}:\gamma)$ $=$ $0$
and
[…]
$\Theta_{\phi,s_{2}}$ $=$ $0$
$\Theta_{\phi,s_{3}}$ $=$ $0.$
On the issue of uniqueness. In the context of a regular parameter $\phi$, the uniqueness of the map $\iota_{\phi}\colon\Pi_{\phi}\to\mathrm{Irr}(S_{\phi})$ is clear. For the $\mathfrak{w}$-generic element in $\Pi_{\phi}$ is sent to the trivial character of $S_{\phi}$, whereas the other element in $\Pi_{\phi}$ is sent to the non-trivial character of $S_{\phi}$.
In the context of the non-regular parameter $\phi$, the left-hand-side of (1.6) depends, in its definition, on the map $\iota_{\phi}\colon\Pi_{\phi}\to\mathrm{Irr}(S_{\phi})$. The endoscopic character identity selects a unique map $\iota_{\phi}:\Pi_{\phi}\to\mathrm{Irr}(S_{\phi})$, see §4.
We discovered a crucial typo in the statement of Theorem 15.2 in [ADSS]; the corrected statement appears in §5.
Acknowledgement. We would like to thank Loren Spice for a valuable exchange of emails, Shaun Stevens for valuable discussions and Tasho Kaletha for sending us constructive criticism.
2. Elliptic tori
2.0.1. Elliptic tori
There are two tori relevant for us: Let $\varepsilon$ be a fixed nonsquare element in $\mathfrak{o}_{F}^{\times}$, then
(2.1) $T^{\varepsilon}:=\left\{\left(\begin{matrix}a&b\cr b\varepsilon&a\end{matrix}\right)\;:\;a,b\in\mathfrak{o}_{F},a^{2}-\varepsilon b^{2}=1\right\}$
and
(2.2) $T^{\varepsilon,\varpi}:=\left\{\left(\begin{matrix}a&b\varpi^{-1}\cr b\varepsilon\varpi&a\end{matrix}\right)\;:\;a,b\in\mathfrak{o}_{F},a^{2}-\varepsilon b^{2}=1\right\}.$
These two tori admit the same splitting field, namely $F(\sqrt{\varepsilon})$ and are therefore stably conjugate, or, equivalently, $\mathrm{GL}_{2}(F)$-conjugate, thanks to the following identity in $\mathrm{GL}_{2}(F)$:
(2.3) $\left(\begin{array}[]{cc}1&0\\ 0&\varpi\end{array}\right)\left(\begin{array}[]{cc}a&b\\ \varepsilon b&a\end{array}\right)\left(\begin{array}[]{cc}1&0\\ 0&\varpi^{-1}\end{array}\right)=\left(\begin{array}[]{cc}a&\varpi^{-1}b\\ \varepsilon\varpi b&a\end{array}\right).$
There is a single stable conjugacy class of unramified elliptic maximal $F$-tori, represented by $T^{\varepsilon}$. It splits into two distinct $G$-conjugacy classes, represented by $T^{\varepsilon}$ and $T^{\varepsilon,\varpi}$.
Let $E^{1}$ denote the group of norm $1$ elements in $E$. We have, following [ADSS, §3.1], canonical isomorphisms of $F$-tori:
$T^{\varepsilon}\to E^{1},\quad\left(\begin{matrix}a&b\cr b\varepsilon&a\end{matrix}\right)\mapsto a+b\sqrt{\varepsilon}$
$T^{\varepsilon,\varpi}\to E^{1},\quad\left(\begin{matrix}a&b\varpi^{-1}\cr b\varepsilon\varpi&a\end{matrix}\right)\mapsto a+b\sqrt{\varepsilon}.$
Given $a+b\sqrt{\varepsilon}\in E^{1}$ we shall, following [ADSS, §3.1] write
$\mathrm{Im}_{\varepsilon}(a+b\sqrt{\varepsilon})=b.$
We have
(2.4) $\mathrm{Im}_{\varepsilon}(\mathbf{g}\cdot\gamma)=\mathrm{Im}_{\varepsilon}(\gamma),$
for every $\gamma\in T^{\varepsilon}$, where
(2.5) $\mathbf{g}:=\left(\begin{array}[]{cc}1&0\\ 0&\varpi\end{array}\right)\in\mathrm{GL}_{2}(F)$
and $\mathbf{g}\cdot\gamma:=\mathbf{g}\gamma\mathbf{g}^{-1}$. The equation (2.3) can be rephrased as the following statement:
(2.6) $\mathbf{g}\cdot T^{\varepsilon}=\mathbf{g}T^{\varepsilon}\mathbf{g}^{-1}=T^{\varepsilon,\varpi}.$
We also note that
$\mathrm{sgn}_{\varepsilon}(x)=(-1)^{\mathrm{v}(x)}$
for all $x\in F$, see (2.1) in [ADSS]. Then we have, using (2.4),
$\mathrm{sgn}_{\varepsilon}(\varpi^{-1}\mathrm{Im}_{\varepsilon}(\mathbf{g}\cdot\gamma))$ $=$ $\mathrm{sgn}_{\varepsilon}(\varpi^{-1}\mathrm{Im}_{\varepsilon}(\gamma))$
$=$ $\mathrm{sgn}_{\varepsilon}(\varpi^{-1})\mathrm{sgn}_{\varepsilon}(\mathrm{Im}_{\varepsilon}(\gamma))$
$=$ $(-1)^{\mathrm{v}(\varpi^{-1})}\mathrm{sgn}_{\varepsilon}(\mathrm{Im}_{\varepsilon}(\gamma))$
$=$ $-\mathrm{sgn}_{\varepsilon}(\mathrm{Im}_{\varepsilon}(\gamma)).$
2.0.2. Weyl discriminant
In [ADSS, Definition 3.7], the Weyl discriminant $D_{G}$ is defined as follows:
$D_{G}\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right)=(a+d)^{2}-4.$
Now
$\gamma=\left(\begin{matrix}a&b\cr b\varepsilon&a\end{matrix}\right)\in T^{\varepsilon}\implies D_{G}(\gamma)=4a^{2}-4=4\varepsilon b^{2}$
so that
$|D_{G}(\gamma)|^{1/2}=|b|.$
2.0.3. The function $f$
We set
(2.8) $f(\gamma):=\frac{\mathrm{sgn}_{\varepsilon}(\mathrm{Im}_{\varepsilon}(\gamma))}{|D_{G}(\gamma)|^{1/2}}.$
Note that
$\gamma\quad\textrm{is regular}\iff b\neq 0.$
In that case, we have
$f(\gamma)$ $=$ $\frac{\mathrm{sgn}_{\varepsilon}(b)}{|b|}$
$=$ $(-1)^{\mathrm{v}(b)}q^{\mathrm{v}(b)}$
$=$ $(-q)^{\mathrm{v}(b)}.$
It is worth noting that $f$ is invariant under the Galois group $\mathrm{Gal}(F(\sqrt{\varepsilon})/F)$. If $c=a+b\sqrt{\varepsilon}$ then we have
(2.9) $f(\overline{c})=f(a-b\sqrt{\varepsilon})=(-q)^{\mathrm{v}(-b)}=(-q)^{\mathrm{v}(b)}=f(c).$
The following Lemma is proved in [ADSS, p.65]. We offer a short proof.
Lemma 2.10.
Let $\gamma$ be a regular semisimple element in the elliptic torus $T^{\varepsilon}$ which is far from the identity. Then we have $f(\gamma)=1$.
Proof.
We recall the pro-unipotent radical $T^{\varepsilon}_{1}$ of $T^{\varepsilon}$, see (1.3). If $\gamma$ is far from the identity, then we have $\gamma\in T^{\varepsilon}\setminus T^{\varepsilon}_{1}$ and so $b\in\mathfrak{o}_{F}\setminus\mathfrak{p}_{F}$. Therefore $\mathrm{v}(b)=0$. ∎
2.0.4. Elliptic torus in $\mathrm{SL}_{2}(\mathfrak{f})$
The quotient $T^{\varepsilon}/T_{1}^{\varepsilon}$ is the unique maximal elliptic $\mathfrak{f}$-torus $\mathbb{T}^{\epsilon}(\mathfrak{f})$ in $\mathrm{SL}_{2}(\mathfrak{f})$:
(2.11) $\mathbb{T}^{\epsilon}(\mathfrak{f})=\left\{\left(\begin{matrix}a&b\\ \epsilon b&a\end{matrix}\right)\,:\,\text{$a,b\in\mathfrak{f}$ and $a^{2}-\epsilon b^{2}=1$}\right\}$
where $\mathfrak{f}$ is the residue field of $F$.
The group $E^{1}$ of norm one elements in $E=F(\sqrt{\varepsilon})$ admits a unique quadratic character $\psi_{0}$ of depth zero which therefore factors through $\mathbb{T}^{\epsilon}$. An explicit formula for $\psi_{0}$ (which we shall not need) is given in [ADSS, §9.2].
3. Depth-zero supercuspidal representations of $\mathrm{SL}_{2}(F)$
Let $I$ denote a line-segment (chamber) in the reduced building (tree) of $\mathrm{SL}_{2}(F)$. The ends of the interval $I$ will be denoted $x_{0}$, $x_{1}$ . The stabilizer of $x_{i}$ will be denoted $G_{x_{i}}$, for $i=0,1$. We have $G_{x_{0}}=\mathrm{SL}_{2}(\mathfrak{o}_{F})$, and $G_{x_{1}}$ is the conjugate of $G_{x_{0}}$ under the element $\mathbf{g}$ of $\mathrm{GL}_{2}(F)$ defined in (2.5). The reductive quotients of $G_{x_{0}}$ and $G_{x_{1}}$ are both isomorphic to $\mathrm{SL}_{2}(\mathfrak{f})$, and $\{G_{x_{0}},G_{x_{1}}\}$ is a set of representatives of the maximal parahoric subgroups of $\mathrm{SL}_{2}(F)$.
Let $\psi$ be a character of $\mathbb{T}^{\epsilon}(\mathfrak{f})$ such that $\psi\neq\psi^{-1}$. The Deligne-Lusztig virtual character $R_{\mathbb{T}^{\epsilon},\psi}^{\mathrm{SL}_{2}}$ is irreducible, and we denote by $|R_{\mathbb{T}^{\epsilon},\psi}|$ its inflation to $G_{x_{0}}$ and $G_{x_{1}}$. We put
(3.1) $\pi^{+}=\pi(\mathbb{T}^{\epsilon},\psi):=c-\mathrm{Ind}_{G_{x_{0}}}^{G}|R_{\mathbb{T}^{\epsilon},\psi}|\quad\text{and}\quad\pi^{-}:=c-\mathrm{Ind}_{G_{x_{1}}}^{G}|R_{\mathbb{T}^{\epsilon},\psi}|.$
The representations $\pi^{+}$ and $\pi^{-}$ are depth-zero regular supercuspidal irreducible representations of $G$ in the terminology of [Ka2].
Let $\psi_{0}$ be the quadratic character of $\mathbb{T}^{\epsilon}(\mathfrak{f})$. The Deligne-Lusztig virtual character $R_{\mathbb{T}^{\epsilon},\psi_{0}}^{\mathrm{SL}_{2}}$ decomposes as
(3.2) $R_{\mathbb{T}^{\epsilon},\psi_{0}}^{\mathrm{SL}_{2}}=-R_{\mathbb{T}^{\epsilon},\psi_{0}}^{+}-R_{\mathbb{T}^{\epsilon},\psi_{0}}^{-},$
where $R_{\mathbb{T}^{\epsilon},\psi_{0}}^{\pm}$ are irreducible cuspidal representations of $\mathrm{SL}_{2}(\mathfrak{f})$ of dimension $\frac{q-1}{2}$ (see [DM, Table 12.1]). We choose the signs as in [ADSS, (9.3)].
We denote by $|R_{\mathbb{T}^{\epsilon},\psi_{0}}^{\pm}|$ the inflations of $R_{\mathbb{T}^{\epsilon},\psi_{0}}^{\pm}$ to $G_{x_{0}}$ and $G_{x_{1}}$. The four compactly induced representations $\mathrm{c\!\!-\!\!Ind}_{G_{x_{i}}}^{G}|R_{\mathbb{T}^{\epsilon},\psi_{0}}^{\pm}|$, $i=0,1$, are depth-zero supercuspidal irreducible representations of $G$, they are called “exceptional” in [ADSS], and are non-singular non-regular in the terminology of [Ka3] (see also [Au2]). We write
(3.3) $\pi^{+}:=\mathrm{c\!\!-\!\!Ind}_{G_{x_{0}}}^{G}|R_{\mathbb{T}^{\epsilon},\psi_{0}}^{+}|\quad\text{and}\quad\pi^{-}:=\mathrm{c\!\!-\!\!Ind}_{G_{x_{1}}}^{G}|R_{\mathbb{T}^{\epsilon},\psi_{0}}^{-}|.$
We define
$(\mathbf{g}\cdot\pi^{+})(x):=\pi^{+}(\mathbf{g}^{-1}x\mathbf{g})\quad\text{and}\quad\mathbf{g}\cdot x:=\mathbf{g}x\mathbf{g}^{-1},$
so that
$(\mathbf{g}\cdot\pi^{+})(x)=\pi^{+}(\mathbf{g}^{-1}\cdot x).$
We set
(3.4) $\pi_{1}=\pi^{+},\quad\pi_{2}=\pi^{-},\quad\pi_{3}=\mathbf{g}\cdot\pi^{+},\quad\pi_{4}=\mathbf{g}\cdot\pi^{-}.$
4. Virtual characters far from the identity: the non-regular parameter
We recall that $G=\mathrm{SL}_{2}(F),\,\widehat{G}=\mathrm{PGL}_{2}(\mathbb{C})=\mathrm{PSL}_{2}(\mathbb{C})$. Let $E/F$ be the (unramified) quadratic extension of $F$ containing $\sqrt{\varepsilon}$ and let $K/F$ be the biquadratic extension of $F$ containing $\sqrt{\varepsilon}$ and $\sqrt{\varpi}$. For the Galois group we have $\mathrm{Gal}(K/F)=\{1,\sigma,\tau,\sigma\tau\}$ with $\sigma\sqrt{\varepsilon}=-\sqrt{\varepsilon}$ and $\tau\sqrt{\varpi}=-\sqrt{\varpi}$.
The Langlands parameter
(4.1) $\phi\colon\mathbf{W}_{F}\to\mathrm{PGL}_{2}(\mathbb{C})$
is now defined by the condition that $\phi$ factors through $\mathrm{Gal}(K/F)$ and satisfies the equations
(4.2) $\phi(\sigma)$ $=$ $s_{2},\quad\phi(\tau)=s_{1}.$
We consider the associated supercuspidal $L$-packet
$\Pi_{\phi}(\mathrm{SL}_{2}(F))=\{\pi_{1},\pi_{2},\pi_{3}.\pi_{4}\},$
and let $\Theta_{j}:=\Theta_{\pi_{j}}$ with $1\leq j\leq 4$.
Theorem 4.3.
Let $\gamma$ be a regular semisimple element in the elliptic torus $T^{\varepsilon}$ which is far from the identity. Then we have the character formulas
$\Theta_{1}(\gamma)$ $=$ $-\psi_{0}(\gamma)$
$\Theta_{2}(\gamma)$ $=$ $-\psi_{0}(\gamma)$
$\Theta_{3}(\gamma)$ $=$ $0$
$\Theta_{4}(\gamma)$ $=$ $0.$
Proof.
We refer to equations ()(*) and ()(**) on p.65 of [ADSS]. It follows immediately that
$\Theta_{\pi^{+}}(\gamma)=\Theta_{\pi^{-}}(\gamma)=-\psi_{0}(\gamma).$
We have, for all $\gamma\in T^{\varepsilon}$,
$\Theta_{\pi_{3}}(\gamma)$ $=$ $\Theta_{\mathbf{g}\cdot\pi^{+}}(\gamma)$
$=$ $\Theta_{\pi^{+}}(\mathbf{g}^{-1}\cdot\gamma)$
$=$ $0$
since $\Theta_{\pi^{+}}$ vanishes on $\mathbf{g}^{-1}\cdot T^{\varepsilon}$ by [ADSS, Theorem 15.1].
Similarly, we have, for all $\gamma\in T^{\varepsilon}$,
$\Theta_{\pi_{4}}(\gamma)$ $=$ $\Theta_{\mathbf{g}\cdot\pi^{-}}(\gamma)$
$=$ $\Theta_{\pi^{-}}(\mathbf{g}^{-1}\cdot\gamma)$
$=$ $0$
since $\Theta_{\pi^{-}}$ vanishes on $\mathbf{g}^{-1}\cdot T^{\varepsilon}$ by [ADSS, Theorem 15.1]. ∎
Consider the $L$-packet of $\mathrm{SL}_{2}(F)$ given by
(4.4) $\Pi_{\phi}(\mathrm{SL}_{2}(F))=\{\pi_{1},\pi_{2},\pi_{3},\pi_{4}\}$
where the representations $\pi_{i}$ are depth-zero supercuspidal irreducible representations of $G$ (see (3.4)). We suppose that $\pi_{1}$ is $\mathfrak{w}$-generic.
Let $\Theta_{1}$, $\Theta_{2}$, $\Theta_{3}$, $\Theta_{4}$ be the distribution characters of $\pi_{1}$, $\pi_{2}$, $\pi_{3}$, $\pi_{4}$.
Given $A\in\mathrm{GL}_{2}(\mathbb{C})$ let $A_{*}$ denote the image of $A$ in $\mathrm{PGL}_{2}(\mathbb{C})$. Let
(4.5) $s_{1}=\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right)_{*}\quad s_{2}=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right)_{*}\quad s_{3}=\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)_{*}.$
We have
$S_{\phi}=\{1,s_{1},s_{2},s_{1}s_{2}\}.$
We will enumerate the characters of $S_{\phi}$ in the following way:
11 $s_{1}$ $s_{2}$ $s_{1}s_{2}$
$\rho_{1}$ 11 11 11 11
$\rho_{2}$ 11 11 1-1 1-1
$\rho_{3}$ 11 1-1 11 1-1
$\rho_{4}$ 11 1-1 1-1 11
Given
(4.6) $\iota\colon\Pi_{\phi}\to\mathrm{Irr}(S_{\phi})$
we attach to $(\phi,s)$ the virtual character
(4.7) $\Theta_{\phi,s}:=\sum_{\pi\in\Pi_{\phi}(\mathrm{SL}_{2}(F))}\iota(\pi)(s)\cdot\Theta_{\pi}.$
The first constraint is that $\iota$ must assign the trivial character $\rho_{1}$ to the $\mathfrak{w}$-generic representation $\pi_{1}$. The second important constraint is the endoscopic character identity. These constraints determine the map $\iota$ uniquely. The defining equation is
(4.8) $\iota(\pi_{j}):=\rho_{j}\quad\text{ with $j=1,2,3,4$.}$
In that case, we have
(4.9) $\Theta_{\phi,s}=\begin{cases}\Theta_{1}+\Theta_{2}+\Theta_{3}+\Theta_{4}&\text{if $s=1$}\\ \Theta_{1}+\Theta_{2}-\Theta_{3}-\Theta_{4}&\text{if $s=s_{1}$}\\ \Theta_{1}-\Theta_{2}+\Theta_{3}-\Theta_{4}.&\text{if $s=s_{2}$}\\ \Theta_{1}-\Theta_{2}-\Theta_{3}+\Theta_{4}&\text{if $s=s_{3}$}\end{cases}.$
Theorem 4.10.
Let $\gamma$ be a regular semisimple element in the elliptic torus $T^{\varepsilon}$ which is far from the identity. Then we have
$\Theta_{\phi,1\hphantom{s}}(\gamma)$ $=$ $-2\psi_{0}(\gamma)$
$\Theta_{\phi,s_{1}}(\gamma)$ $=$ $-2\psi_{0}(\gamma)$
$\Theta_{\phi,s_{2}}(\gamma)$ $=$ $0$
$\Theta_{\phi,s_{3}}(\gamma)$ $=$ $0.$
Proof.
This follows readily from Theorem 4.3. ∎
5. Virtual characters near the identity: the non-regular parameter
The sign character $\mathrm{sgn}_{\varepsilon}\colon F^{\times}\to\{\pm 1\}$ is determined by the short exact sequence
$1\to{\mathrm{N}}_{E/F}E^{\times}\to F^{\times}\to\mathrm{Gal}(E/F)\to 1.$
The notation $\mathrm{Im}_{\varepsilon}$ has already been defined in §3.1. We recall the definition of $f(\gamma)$ from (2.8).
Theorem 5.1.
Let $\gamma$ be a regular semisimple element in the elliptic torus $T^{\varepsilon}$ which is near the identity. Then we have
$\Theta_{\phi,1}(\gamma)$ $=$ $-2$
$\Theta_{\phi,s_{1}}(\gamma)$ $=$ $-2f(\gamma)$
$\Theta_{\phi,s_{2}}(\gamma)$ $=$ $0$
$\Theta_{\phi,s_{3}}(\gamma)$ $=$ $0$
Proof.
According to Theorem 15.2 in [ADSS], we have 111we have made a surely necessary correction to [ADSS, Theorem 15.2], deleting the $\pm$ inside the curly bracket
$\Theta_{\pi^{\pm}}(\gamma)=\frac{1}{2}\left\{H(\Lambda^{\prime},F_{\theta^{\prime}})\frac{\mathrm{sgn}_{\theta^{\prime}}(\eta^{-1}\mathrm{Im}_{\theta^{\prime}}(\gamma))}{|D_{G}(\gamma)|^{1/2}}-1\right\}$
With $\gamma\in T^{\varepsilon},\theta^{\prime}=\varepsilon,\eta=1$ we have
$\Theta_{1}(\gamma)$ $=$ $\Theta_{2}(\gamma)$
$=$ $(-f(\gamma)-1)/2$
With $\gamma\in T^{\varepsilon},\mathbf{g}\cdot\gamma\in T^{\varepsilon,\varpi},\theta^{\prime}=\varepsilon,\eta=\varpi$ we have
$\Theta_{4}(\gamma)$ $=$ $\Theta_{3}(\gamma)$
$=$ $\Theta_{1}(\mathbf{g}\cdot\gamma)$
$=$ $(f(\gamma)-1)/2$
and the result follows immediately. ∎
6. Virtual characters: regular parameters
We recall from (3.1) the definition of the depth-zero supercuspidal representation
$\pi^{+}:=\pi(T^{\varepsilon},\psi).$
The $L$-packet is then
$\Pi_{\phi}=\{\pi^{+},\pi^{-}\}$
with $\pi^{-}=\mathbf{g}\cdot\pi^{+}$.
Theorem 6.1.
Let $s$ be the non-trivial element in $S_{\phi}$. Let $\gamma$ be a regular semisimple element in the elliptic torus $T^{\varepsilon}$. Then we have
[…]
Proof.
First case: $\gamma$ is far from the identity. The equation
$\Theta_{\pi^{+}}(\gamma)=-\psi(\gamma)-\psi(1/\gamma)$
is a re-statement of Theorem 14.14 in [ADSS]. We have simplified this statement by using Lemma 2.10 and [ADSS, Lemma 4.2]. According to [ADSS, Theorem 14.14] we have
$\Theta_{\pi^{-}}(\gamma)=0.$
We recall that $S_{\phi}=\{1,s\}=\mathbb{Z}/2\mathbb{Z}$. Denote the trivial character of $S_{\phi}$ by $\rho_{0}$ and the non-trivial character by $\rho_{1}$. Assume that $\pi$ is $\mathfrak{w}$-generic so that the parametrization of the $L$-packet is
$\iota\colon\Pi_{\phi}\to\mathrm{Irr}(S_{\phi}),\quad\quad\pi^{+}\mapsto\rho_{0},\quad\pi^{-}\mapsto\rho_{1}.$
In that case, we have
$\Theta_{\phi,s}(\gamma)$ $=$ $\sum_{\pi\in\Pi_{\phi}(\mathrm{SL}_{2}(F))}\mathrm{trace}(\iota(\pi))(s)\cdot\Theta_{\pi}$
$=$ $\rho_{0}(s)\Theta_{\pi}(\gamma)+\rho_{1}(s)\Theta_{\pi^{-}}(\gamma)$
$=$ $\Theta_{\pi^{+}}(\gamma)$
$=$ $-\psi(\gamma)-\psi(1/\gamma).$
Second case: $\gamma$ is near the identity. In the terminology of [ADSS, Definition 14.1], we are in the unramified case. According to [ADSS, Theorem 14.20] we have
$\Theta_{\pi^{+}}(\gamma)=c_{0}(\pi)-f(\gamma)$
where, by [ADSS, Definition 14.17], the constant term $c_{0}(\pi)$ is
$c_{0}(\pi)=-1$
and so
$\Theta_{\pi^{+}}(\gamma)=-1-f(\gamma).$
Note that $\gamma\in T^{\varepsilon}\implies\mathbf{g}\cdot\gamma\in T^{\varepsilon,\varpi}$. With $\gamma\in T_{1}^{\varepsilon}$ we have
$\Theta_{\pi^{+}}(\mathbf{g}\cdot\gamma)$ $=$ $c_{0}(\pi)-\frac{\mathrm{sgn}_{\varepsilon}(\varpi^{-1}\mathrm{Im}_{\varepsilon}(\mathbf{g}\cdot\gamma))}{D^{G}(\mathbf{g}\cdot\gamma)}$
by Theorem 14.20 in [ADSS]. Using 2.0.1, we obtain
$\Theta_{\pi^{+}}(\mathbf{g}\cdot\gamma)=c_{0}(\pi)+f(\gamma)$
With $\gamma\in T^{\varepsilon}$ we also have
$\Theta_{\pi^{-}}(\gamma)$ $=$ $\Theta_{\pi^{+}}(\mathbf{g}\cdot\gamma)$
and so
$\Theta_{\pi^{-}}(\gamma)=-1+f(\gamma).$
The virtual character is given by
$\Theta_{\phi,s}(\gamma)$ $=$ $\rho_{0}(s)\Theta_{\pi^{+}}(\gamma)+\rho_{1}(s)\Theta_{\pi^{-}}(\gamma)$
$=$ $\Theta_{\pi^{+}}(\gamma)-\Theta_{\pi^{-}}(\gamma)$
$=$ $-1-f(\gamma)-(-1+f(\gamma))$
$=$ $-2f(\gamma).$
7. Endoscopy: the non-regular parameter
We recall from §4 the definition of the non-regular parameter $\phi$. The defining equations are
$\phi(\sigma)=s_{2},\quad\quad\phi(\tau)=s_{1}.$
We have
Theorem 7.1.
Let $E=F(\sqrt{\varepsilon})$ and let $H$ denote the norm $1$ subgroup of $E^{\times}$. Let $\psi_{0}$ denote the quadratic character of $H$. The non-regular parameter $\phi$ factrorizes as follows:
$\begin{CD}\phi:\mathbf{W}_{F}@>{\phi^{H}}>{}>\widehat{H}\rtimes\mathbf{W}_{F}@>{{}^{L}\eta}>{}>\widehat{G}\times\mathbf{W}_{F}\end{CD}$
and the stable character is given by
$\mathscr{S}\Theta_{\phi^{H}}=\psi_{0}.$
Proof.
We consider the identity component of the centralizer of $s_{1}$ in $\widehat{G}$. This is a complex torus of dimension $1$. Following [T, §6.2] we will set $\widehat{H}=\mathrm{Cent}(s_{1},\widehat{G})^{0}$ and consider $\mathcal{H}=\widehat{H}\cdot\phi(\mathbf{W}_{F})$ and the inclusion
$^{L}\eta:\mathcal{H}\to{}^{L}G.$
Note that
$\widehat{H}\cdot\phi(\mathbf{W}_{F})=\widehat{H}\cdot\phi(\sigma)$
since $\phi(\tau)\in\widehat{H}$. The maximal compact subgroup of $\mathrm{PGL}_{2}(\mathbb{C})$ is $\mathrm{SO}_{3}(\mathbb{R})$ and $s_{1},s_{2}$ may be viewed as rotations of order $2$ about orthogonal axes. Then any $z\in\widehat{H}$ is a rotation about the axis of $s_{1}$ and the conjugate of $z$ by the rotation $s_{2}$ is precisely the inverse of the rotation $z$.
So $\phi(\sigma)$ acts on $\widehat{H}$ as inversion and this determines an action of $\mathrm{Gal}(E/F)$ by inversion on $\widehat{H}$. Let $H$ be the quasi-split group defined over $F$ that is dual to $\widehat{H}$ and whose rational structure is determined by $\mathrm{Gal}(E/F)\to\mathrm{Out}(H)$ as above. The group of norm $1$ elements in $E$ admits a canonical action of $\mathrm{Gal}(E/F)$ by inversion and so qualifies as the endoscopic group $H$.
The parameter $\phi$ factors through $^{L}H$:
$\begin{CD}\phi:\mathbf{W}_{F}@>{\phi^{H}}>{}>\widehat{H}\rtimes\mathbf{W}_{F}@>{{}^{L}\eta}>{}>\widehat{G}\times\mathbf{W}_{F}.\end{CD}$
It is worth noting that the centralizer of $s_{1}$ in $\widehat{G}$ is precisely $^{L}H$.
The maps $\phi^{H}$ and $^{L}\eta$ are given explicitly by
$\phi^{H}(\sigma)=1\rtimes\sigma,\quad\quad\phi^{H}(\tau)=s_{1}\rtimes\tau$
and
$^{L}\eta(1\rtimes\sigma)=s_{2}\times\sigma,\quad{}^{L}\eta(s_{1}\rtimes\tau)=s_{1}\times\tau$
We now have the extended endoscopic triple
$\mathfrak{e}_{1}:=({}^{L}H,s_{1},{}^{L}\eta)$
In the notation of [FKS, §4,4] we have $(S,\theta)=(H,\psi_{0})$. Then $\phi^{H}$ is the $L$-parameter of the quadratic character $\psi_{0}$. Therefore, we have
$\Theta_{\phi^{H}}=\psi_{0}$
and, for the stable character,
(7.2) $\mathscr{S}\Theta_{\phi^{H}}=\psi_{0}.$
Lemma 7.3.
Let $\gamma$ be a regular semisimple element in $T^{\varepsilon}$. Then we have
$\Delta[\mathfrak{w},\mathfrak{e}_{1}](\gamma,\gamma)$ $=$ $-f(\gamma)$
$\delta\in H_{2}\implies\Delta[\mathfrak{w},\mathfrak{e}_{2}](\delta,\gamma)$ $=$ $0$
$\delta\in H_{3}\implies\Delta[\mathfrak{w},\mathfrak{e}_{3}](\delta,\gamma)$ $=$ $0$
with $\mathfrak{e}_{i}=\mathfrak{e}(s_{i})$ for $i=1.2.3$.
Proof.
Underlying the map $\widehat{\eta}\colon\widehat{H}\to\widehat{G}$ we have a canonical pair of isomorphisms of $F$-tori:
$T^{\varepsilon}$ $\to$ $H$
$\left(\begin{array}[]{cc}a&b\\ \varepsilon b&a\end{array}\right)$ $\mapsto$ $a+b\sqrt{\varepsilon}$
and
$\left(\begin{array}[]{cc}a&b\\ \varepsilon b&a\end{array}\right)$ $\mapsto$ $a-b\sqrt{\varepsilon}.$
These isomorphisms are admissible in the sense of [Ka2, §3]. We therefore have two related pairs:
(7.6) $(\gamma,\gamma)\quad\textrm{and}\quad(1/\gamma,\gamma).$
We will now quote the Kaletha formula for the absolute transfer factor [Ka4, Example 3.6.9]. In our notation, this is
(7.7) $\epsilon(1/2,\mathbb{C},\psi)\cdot\kappa_{E/F}\left(\frac{c-\overline{c}}{2\eta}\right)|c-\overline{c}|_{F}.$
We recall that
$D_{G}(\gamma)=|c-\overline{c}|_{F}=|b|.$
In order to match the normalization of transfer factors in the recent article [AK], we switch the position of the Weyl discriminant $|c-\overline{c}|_{F}$ from numerator to denominator. The transfer factor defined in [AK], for which there is a clean formulation of the endoscopic character identity, is the transfer factor in [Ka1, p.163] divided by $D_{G}(\gamma)^{2}$. This leads to the equation
(7.8) $\Delta[\mathfrak{w},\mathfrak{e}_{1}](c,c)=\epsilon(1/2,\mathbb{C},\psi)\cdot\kappa_{E/F}\left(\frac{c-\overline{c}}{2\eta}\right)\cdot|c-\overline{c}|_{F}^{-1}.$
It follows that
$\Delta[\mathfrak{w},\mathfrak{e}_{1}](c,c)$ $=$ $\epsilon(1/2,\mathbb{C},\psi)\cdot\mathrm{sgn}_{\varepsilon}(b)\cdot|b|^{-1}.$
Local class field theory allows us to replace the $\mathbf{W}_{F}$-module $\mathbb{C}$ with the sign character $\mathrm{sgn}_{\varepsilon}$. Denote this character by $\chi$. We infer that
$\epsilon(1/2,\mathbb{C},\psi)=\epsilon(\chi,1/2,\psi)$
the Tate local constant of $\chi$ as in [BH, §23.4]. Note that $\Lambda$ has depth zero” in [ADSS] is the same as $\psi$ has level one” in [BH]. We now apply the Proposition in [BH, §23.5] and we obtain
$\epsilon(1/2,\mathbb{C},\psi)$ $=$ $\epsilon(\chi,1/2,\psi)$
$=$ $\chi(\varpi)^{-1}$
$=$ $-1.$
Putting all this together, we conclude that
(7.9) $\Delta[\mathfrak{w},\mathfrak{e}_{1}](\gamma,\gamma)=-f(\gamma).$
According to (2.9), $f$ is invariant under the action of $\mathrm{Gal}(E/F)$, from which we infer that
$\Delta[\mathfrak{w},\mathfrak{e}_{1}](1/\gamma,\gamma)=-f(\gamma).$
Set $\widehat{H_{2}}=\mathrm{Cent}(s_{2},\widehat{G})^{0}$. Then
$\widehat{H_{2}}\cdot\phi(\mathbf{W}_{F})=\widehat{H_{2}}\cdot\phi(\tau)$
and $\phi(\tau)$ acts on $\widehat{H_{2}}$ by inversion. Therefore, $H_{2}$ is secured as the group of norm $1$ elements in $E_{2}=F(\sqrt{\varpi})$. We have the extended endoscopic triple
$\mathfrak{e}_{2}=(s_{2},H_{2},{}^{L}\eta_{2}).$
Set $\widehat{H_{3}}=\mathrm{Cent}(s_{3},\widehat{G})^{0}$. Then
$\widehat{H_{3}}\cdot\phi(\mathbf{W}_{F})=\widehat{H_{3}}\cdot\phi(\sigma)\phi(\tau)$
and $\phi(\sigma)$, $\phi(\tau)$ separately act on $\widehat{H_{3}}$ by inversion. Therefore, $H_{3}$ is secured as the group of norm $1$ elements in $E_{3}=F(\sqrt{\varepsilon\varpi})$. We have the extended endoscopic triple
$\mathfrak{e}_{3}=(s_{3},\mathrm{H}_{3},{}^{L}\eta_{3}).$
At this point, write $H=H_{1}$ and $E=E_{1}$. Note that $H_{1}$ (resp. $H_{2},H_{3}$) is the endoscopic group $U(1)$ split over $E_{1}$ (resp. $E_{2},E_{3}$), as in [CFM+, §11.1.5].
There is no admissible isomorphism from $T^{\varepsilon}$ to $H_{2}$. If $\delta\in H_{2}$ and $\gamma\in T^{\varepsilon}$ then $\gamma$ and $\delta$ are not related and so
$\Delta[\mathfrak{w},\mathfrak{e}_{2}](\delta,\gamma)=0.$
Similarly, there is no admissible isomorphism from $T^{\varepsilon}$ to $H_{3}$. If $\delta\in H_{3}$ and $\gamma\in T^{\varepsilon}$ then $\gamma$ and $\delta$ are not related and so
$\Delta[\mathfrak{w},\mathfrak{e}_{3}](\delta,\gamma)=0.$
Theorem 7.10.
Let $\phi$ be the non-regular supercuspidal Langlands parameter for $\mathrm{SL}_{2}$. If $\gamma\in T^{\varepsilon}$ then we have
$\mathscr{E}(s_{1}:\gamma)$ $=$ $-2f(\gamma)\psi_{0}(\gamma)$
$\mathscr{E}(s_{2}:\gamma)$ $=$ $0$
$\mathscr{E}(s_{3}:\gamma)$ $=$ $0.$
Proof.
We have
$\mathscr{E}(s_{1}:\gamma)$ $=$ $\sum_{\delta\in H(F)}\Delta[\mathfrak{w},\mathfrak{e}_{1}](\delta,\gamma)\,\mathscr{S}\Theta_{\phi^{H}}(\delta)$
[…]
$=$ $\Delta[\mathfrak{w},\mathfrak{e}_{1}](\gamma,\gamma)\psi_{0}(\delta)+\Delta([\mathfrak{w},\mathfrak{e}_{1}](1/\gamma,\gamma))\psi_{0}(1/\gamma)$
[…]
$=$ $-2f(\gamma)\psi_{0}(\gamma)\quad\textrm{since}\quad\psi_{0}=\psi_{0}^{-1}.$
We also have
$\mathscr{E}(s_{2}:\gamma)$ $=$ $\sum_{\delta\in H(F)}\Delta[\mathfrak{w},\mathfrak{e}_{2}](\delta,\gamma)\,\mathscr{S}\Theta_{\phi^{H}}(\delta)$
[…]
and
$\mathscr{E}(s_{3}:\gamma)$ $=$ $\sum_{\delta\in H(F)}\Delta[\mathfrak{w},\mathfrak{e}_{3}](\delta,\gamma)\,\mathscr{S}\Theta_{\phi^{H}}(\delta)$
[…]
Far from the identity, the right-hand-side of Theorem 7.10 is
$2\psi_{0}(\gamma)$
since $f(\gamma)=1$ by Lemma 2.10. Near the identity, the right-hand-side of Theorem 7.10 is
$2f(\gamma)$
since $\gamma$ and $1/\gamma$ are both near the identity and $\psi_{0}$ has depth zero. This mirrors the calculations in [ADSS]: in that paper, the character formulas are done separately, according as $\gamma$ is far from the identity or near the identity.
8. Endoscopy: regular parameters
We recall the elements $s_{1}$, $s_{2}$ and $s_{3}$ in $\mathrm{PGL}_{2}(\mathbb{C})$ defined in (4.5). Recall the unramified quadratic extension $E/F$ and let $\sigma$ be the generator of $\mathrm{Gal}(E/F)$. Let $\xi$ be a character of $E^{\times}$ such that
$\xi\neq\xi^{\sigma}\quad\textrm{and}\quad\psi=\xi|_{E^{1}}.$
Then $(E/F,\xi)$ is an admissible pair in the sense of [BH, §18]. Let
$\rho_{\xi}=\mathrm{Ind}_{E}^{F}\,\xi$
be the induced $\mathbf{W}_{F}$-module. This is a $2$-dimensional imprimitive $\mathbf{W}_{F}$-module. We describe this module explicitly.
We note that $\mathbf{W}_{F}/\mathbf{W}_{E}\simeq\mathrm{Gal}(E/F)$. Coset representatives are $\{1,\sigma\}$.
As a vector space, the induced module is $\mathbb{C}\oplus\mathbb{C}$ with module structure as follows:
  • $w\in\mathbf{W}_{E}$ acts on the first summand $\mathbb{C}$ as $\xi(w)$
  • $w\in\mathbf{W}_{E}$ acts on the second summand $C$ as $\xi^{\sigma}(w)$
  • $\sigma$ sends $(z_{1},z_{2})$ to $(z_{2},z_{1})$.
This $\mathbf{W}_{F}$-module determines a homomorphism
$\rho_{\xi}\colon\mathbf{W}_{F}$ $\to$ $\mathrm{GL}_{2}(\mathbb{C})$
$w$ $\mapsto$ $\left(\begin{array}[]{cc}\xi(w)&0\\ 0&\xi^{\sigma}(w)\end{array}\right)\quad\textrm{if}\quad w\in\mathbf{W}_{E}$
$\sigma$ $\mapsto$ $\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right).$
Then $\phi$ is the composition
(8.3) $\begin{CD}\phi\colon W_{F}@>{\rho_{\xi}}>{}>\mathrm{GL}_{2}(\mathbb{C})@ >>>\mathrm{PGL}_{2}(\mathbb{C}).\end{CD}$
The centralizer of $\mathrm{im}\,\phi$ in $\mathrm{PGL}_{2}(\mathbb{C})$ is $\{1,s_{1}\}$ and so we have
$S_{\phi}=\mathbb{Z}/2\mathbb{Z}.$
We consider the representations $\pi:=\pi^{+}=\pi(\mathbb{T}^{\epsilon},\psi)$ and $\pi^{-}$ defined in (3.1).
Lemma 8.4.
We have
$\Pi_{\phi}(\mathrm{SL}_{2}(F))=\{\pi,\pi^{-}\}.$
Proof.
Given the admissible pair $(E/F,\xi)$ the construction of the irreducible supercuspidal representation $\pi_{\xi}$ is described in [BH, 19.2]. Then we have the Langlands correspondence for $\mathrm{GL}_{2}(F)$ in [BH, 34.4]:
$\rho_{\xi}\mapsto\pi_{\Delta_{\xi}\xi}$
We note that, since $E/F$ is an unramified quadratic extension, the auxiliary character $\Delta_{\xi}$ is unramified of order $2$ by the Definition in [BH, 34.4]. Also $\kappa_{E/F}$ is unramified of order $2$ by the Proposition in [BH, 34.3]. Therefore we have $\Delta_{\xi}=\kappa_{E/F}$ and twisting by $\Delta_{\xi}$ has no effect on $\rho_{\xi}$ by the Lemma in [BH, 34.1]. So the Langlands correspondence is given by
$\rho_{\xi}\mapsto\pi_{\xi}.$
Now we refer to the definition of $\pi^{+}$ and $\pi^{-}$ in (3.1). ∎
Let $\rho_{0}$ denote the trivial character of $S_{\phi}$ and let $\rho_{1}$ denote the non-trivial character of $S_{\phi}$. We will assume that $\pi^{+}$ is $\mathfrak{w}$-generic. Then we have a bijection
$\iota\colon\Pi_{\phi}(\mathrm{SL}_{2}(F))\to\mathrm{Irr}(S_{\phi}),\quad\quad\iota(\pi^{+})=\rho_{0},\quad\iota(\pi^{-})=\rho_{1}.$
Let $s=s_{1}$. Let $\widehat{H}$ denote the identity component of the centralizer of $s$ is $\mathrm{PGL}_{2}(\mathbb{C})$. Define
$\mathcal{H}:=\widehat{H}\cdot\phi(W_{F})$
and let $^{L}\eta:\mathcal{H}\to\mathrm{PGL}_{2}(\mathbb{C})$ denote inclusion. We recall that $E$ is the unramified quadratic extension of $F$. Noting that $\phi(\mathbf{W}_{E})\subset\widehat{H}$, we have
$\mathcal{H}=\widehat{H}\cdot\phi(\sigma).$
This determines an action of $\Gamma_{E/F}$ by conjugation on $\widehat{H}$. This creates the extended endoscopic triple
$\mathfrak{e}=(s,H,{}^{L}\eta).$
We note that this extended endoscopic triple is identical to the one constructed for the non-regular parameter. In the two cases (regular parameter, non-regular parameter) the endoscopic data are the same, but the little parameters $\phi^{H}$ are different.
Theorem 8.5.
Let $\gamma$ be a regular semisimple element in $T^{\varepsilon}$. We have
$\sum_{\delta\in H(F)}\Delta[\mathfrak{w},\mathfrak{e}](\delta,\gamma)\mathscr{S}\Theta_{\phi^{H}}(\delta)=-f(\gamma)(\psi(\gamma)+\psi(1/\gamma)).$
Proof.
We have
$\begin{CD}\phi\colon\mathbf{W}_{F}@>{\phi^{H}}>{}>\mathcal{H}@>{{}^{L}\eta}>{}>\mathrm{PGL}_{2}(\mathbb{C})\end{CD}$
where the little map $\phi^{H}$ is the $L$-parameter of $\psi$ as in [Ka2, p.1148]. Since $\phi^{H}$ is the $L$-parameter of $\psi$ we have $\Theta_{\phi^{H}}=\psi$ and therefore
(8.6) $\mathscr{S}\Theta_{\phi^{H}}=\psi.$
The transfer factor is given by the formula (7), namely
(8.7) $\Delta[\mathfrak{w},\mathfrak{e}](\gamma,\gamma)=\Delta[\mathfrak{w},\mathfrak{e}](1/\gamma,\gamma)=-f(\gamma).$
Finally, we have
[…]
as required. ∎
9. Stability across the inner forms of $\mathrm{SL}_{2}$
Let $\phi$ denote the non-regular parameter of $\mathrm{SL}_{2}$. The finite group $S_{\phi}$ admits a pull-back from the adjoint group $\mathrm{PGL}_{2}(\mathbb{C})$ to its simply-connected cover $\mathrm{SL}_{2}(\mathbb{C})$. The pull-back group $\widetilde{S}_{\phi}$ is the quaternion group $Q$, which admits four one-dimensional representations $\rho_{1},\rho_{2},\rho_{3},\rho_{4}$ and one two-dimensional representation $\rho_{5}$. We have a bijective map
$\mathrm{Irr}(\widetilde{S}_{\phi})\simeq\{\pi_{1},\pi_{2},\pi_{3},\pi_{4},\pi_{5}\}.$
The $L$-packet $\Pi_{\phi}$ is therefore part of a compound packet, comprising five representations, where $\pi_{5}$ is a supercuspidal representation of the inner form $\mathrm{SL}_{1}(D)$. Let $\Theta_{5}$ denote the Harish-Chandra character of $\pi_{5}$.
Theorem 9.1.
The character of the supercuspidal representation $\pi_{5}$ is given by
$\Theta_{5}(\delta)=\psi_{0}(\gamma)$
with $\delta\in\mathrm{SL}_{1}(D)$ and $\gamma\in\mathrm{SL}_{2}(F)$ whenever $\gamma,\delta$ are related.
Proof.
We will consider $\mathrm{GL}_{2}(F)$. Now $\mathrm{GL}_{2}(F)$ admits one inner twist, namely $\mathrm{GL}_{1}(D)$ where the division algebra $D$ has the Brauer invariant
$\mathrm{inv}_{F}(D)=1/2\in\mathbb{Q}/\mathbb{Z}.$
Let $\pi_{0}$ be the depth-zero supercuspidal representation of $\mathrm{GL}_{2}(F)$ for which
$\pi_{0}|_{\mathrm{SL}_{2}(F)}=\pi_{1}\oplus\pi_{2}\oplus\pi_{3}\oplus\pi_{4}.$
Let $\mathrm{JL}$ denote the Jacquet-Langlands correspondence and let $\pi:=\mathrm{JL}^{-1}\pi_{0}$. The local Langlands correspondence for $\mathrm{GL}_{1}(D)$ is determined by the composition
$\begin{CD}\mathrm{Irr}^{2}(\mathrm{GL}_{1}(D))@>{\mathrm{JL}}>{}>\mathrm{Irr}^{2}(\mathrm{GL}_{2}(F))@>{\mathrm{{LLC}}}>{}>\Phi(\mathrm{GL}_{2}).\end{CD}$
The two elements in the set
$\{\pi,\mathrm{JL}(\pi)\}$
therefore admit the same $L$-parameter $\phi$, and comprise a compound packet. The corresponding character identity is
(9.2) $\Theta_{\pi}(\delta)=-\Theta_{\mathrm{JL}(\pi)}(\gamma)$
with $\delta\in\mathrm{GL}_{1}(D)$ and $\gamma\in\mathrm{GL}_{2}(F)$ whenever $\gamma,\delta$ are related. This character identity remains true when we restrict to $\mathrm{SL}_{1}(D)$ and $\mathrm{SL}_{2}(F)$ respectively.
Let $Q$ denote the classical quaternion group. According to [HS, Lemma 12.6], we have, as a representation space for $Q\times\mathrm{SL}_{2}(F)$,
$V_{\pi(\phi)}=\bigoplus_{1\leq j\leq 4}\rho_{j}\otimes\pi_{j}.$
As a representation space for $\mathrm{SL}_{2}(F)$, we therefore have
$V_{\pi(\phi)}=\pi_{1}\oplus\pi_{2}\oplus\pi_{3}\oplus\pi_{4}.$
According to [HS, Lemma 12.6], we have, as a representation space for $Q\times\mathrm{SL}_{1}(D)$,
$V_{\pi(\phi)}=\rho_{5}\otimes\pi_{5}.$
As a representation space for $\mathrm{SL}_{1}(D)$, we therefore have
$V_{\pi(\phi)}$ $=$ $\pi_{5}\oplus\pi_{5}.$
The character identity now follows from (9.2) and we obtain
(9.3) $2\Theta_{5}(\delta)=-(\Theta_{1}+\Theta_{2}+\Theta_{3}+\Theta_{4})(\gamma)$
whenever $\gamma,\delta$ are related with $\gamma\in\mathrm{SL}_{2}(F)_{\mathrm{reg}},\delta\in\mathrm{SL}_{1}(D)_{\mathrm{reg}}$.
We also have
(9.4) $\Theta_{1}+\Theta_{2}+\Theta_{3}+\Theta_{4}$ $=$ $-2\psi_{0}.$
Far from the identity, this follows from Theorem 4.10; near the identity, this follows from Theorem 5.1 since $\psi_{0}$ has depth zero.
By comparing (9.3) with (9.4) we conclude that the character of the supercuspidal representation $\pi_{5}$ is given by
$\Theta_{5}(\delta)=\psi_{0}(\gamma)$
with $\delta\in\mathrm{SL}_{1}(D)$ and $\gamma\in\mathrm{SL}_{2}(F)$ whenever $\gamma,\delta$ are related. ∎
The following corollary is a special case of Theorem 4.4.4(1) in [FKS]. We give an elementary proof.
Corollary 9.5.
The virtual character $\mathscr{S}\Theta_{\phi,*}$ is stable across the two inner forms of $\mathrm{SL}_{2}$:
$\mathscr{S}\Theta_{\phi,0}(\gamma)=\mathscr{S}\Theta_{\phi,1}(\delta)$
for all stably conjugate strongly regular elements $\gamma\in\mathrm{SL}_{2}(F),\delta\in\mathrm{SL}_{1}(D)$.
Proof.
We have
$\mathscr{S}\Theta_{\phi,0}$ $=$ $e(\mathrm{SL}_{2}(F))\cdot(\Theta_{1}+\Theta_{2}+\Theta_{3}+\Theta_{4})$
$=$ $\Theta_{1}+\Theta_{2}+\Theta_{3}+\Theta_{4}$
$=$ $-2\Theta_{5}$
$=$ $e(\mathrm{SL}_{1}(D))\cdot 2\Theta_{5}$
$=$ $\mathscr{S}\Theta_{\phi,1}$
as required. ∎
[…]